### v1

# Less is More: Unlocking Specialization of Time Series Foundation Models via Structured Pruning

Abstract
Scaling laws motivate the development of Time Series Foundation Models (TSFMs) that pre-train vast parameters and achieve remarkable zero-shot forecasting performance. Surprisingly, even after fine-tuning, TSFMs cannot consistently outperform smaller, specialized models trained on full-shot downstream data. A key question is how to realize effective adaptation of TSFMs for a target forecasting task. Through empirical studies on various TSFMs, the pre-trained models often exhibit inherent sparsity and redundancy in computation, suggesting that TSFMs have learned to activate task-relevant network substructures to accommodate diverse forecasting tasks. To preserve this valuable prior knowledge, we propose a structured pruning method to regularize the subsequent fine-tuning process by focusing it on a more relevant and compact parameter space. Extensive experiments on seven TSFMs and six benchmarks demonstrate that fine-tuning a smaller, pruned TSFM significantly improves forecasting performance compared to fine-tuning original models. This “prune-then-finetune” paradigm often enables TSFMs to achieve state-of-the-art performance and surpass strong specialized baselines.
1 Introduction
Time Series Foundation Models (TSFMs) are large forecasting models pre-trained on massive datasets. Recent researches on TSFMs [11, 25, 12, 2, 6, 28, 21] have demonstrated their impressive zero-shot forecasting capabilities across various domains without being specifically trained for those particular prediction tasks. Time series forecasting has a distinct advantage compared to other machine learning tasks. That is, new data naturally accumulates over time, providing fresh training samples without manual labeling. In many industrial sectors such as traffic monitoring, energy consumption, weather prediction, and financial analysis, organizations continuously collect extensive time series data. This data serves two purposes: it provides a long historical context for zero-shot forecasting and abundant samples for model training. Given this data availability, it becomes crucial to evaluate whether generic TSFMs offer real advantages over specialized models trained on task-specific datasets.
Our performance evaluations in Fig. 1 demonstrate that even state-of-the-art pretrained TSFMs operating in zero-shot mode frequently underperform compared to specialized models like PatchTST [23]. This performance gap primarily stems from statistical differences between the diverse pre-training data and the specific patterns in downstream time series. Although fine-tuning TSFM parameters on target datasets improves prediction accuracy, our experiments show that even these adapted TSFMs struggle to consistently outperform PatchTST across various benchmarks. This raises a crucial research question: how can we more effectively leverage these powerful TSFMs to achieve superior performance in specific forecasting applications?
To answer the question, we investigate the internal characteristics of TSFMs that help explain their behaviors. Our analysis reveals that the pre-trained TSFMs often exhibit inherent sparsity in hidden states and redundancy in parameters. For instance, up to 50% of attention heads yield negligible outputs throughout a downstream task, and up to 60% of intermediate channels in feed-forward networks are never activated. These sparsity patterns vary across datasets, suggesting that TSFMs activate different subnetworks depending on the specific task. While conventional fine-tuning retains the original model architecture, it may disrupt these task-relevant activation patterns by modifying all parameters indiscriminately. This approach risks overfitting and fails to capitalize on the valuable task-specific substructures identified during pre-training.
In this paper, we propose decomposing TSFM deployment into two stages: architectural specialization through structured pruning [16, 4] and weight specialization through fine-tuning. While structure pruning by removing redundant components like channels or attention heads is commonly used to compress large models for lower latency and memory usage [22, 27], our primary objective is to facilitate fine-tuning in a narrowed parameter space for stable performance improvement. Specifically, for each downstream task, we analyze channel importance across all linear transformation layers of a TSFM using a loss-guided metric [22] to identify and remove unimportant model parameters. This results in a sparse model whose remaining parameters are fine-tuned on the downstream data.
We conduct extensive experiments on various TSFMs and six popular time series forecasting benchmarks. The results demonstrate that fine-tuning a pruned TSFM yields better forecasting performance than fine-tuning the original model in most cases, achieving up to 22% error reduction. In particular, when selecting the best TSFM on a specific benchmark, our prune-then-finetune approach increases the win rate against PatchTST from 90% to 100%. Moreover, a pruned model achieves superior zero-shot performance when forecasting other time series within the same domain, indicating that the identified model substructure transfers effectively across similar datasets. These findings highlight the essential role of architectural specialization for TSFMs and establish the prune-then-finetune paradigm as a promising strategy for effective TSFM deployment. The major contributions of the paper are summarized in the following.
Identifying the TSFM specialization bottleneck: we provide comprehensive and fair benchmark comparisons showing that even fine-tuned TSFMs struggle to outperform specialized baselines in full-shot forecasting scenarios, highlighting TSFM specialization as a critical unresolved challenge.
Uncovering inherent sparsity in TSFMs: our empirical analysis reveals that pre-trained TSFMs, regardless of whether they utilize sparse Mixture-of-Experts architectures, activate only task-relevant parameter subsets during downstream inference. We argue that these specialized architectural components represent valuable prior knowledge acquired during pre-training and should be preserved when forecasting time series from specific domains.
Pioneering pruning for TSFMs: we propose a novel prune-then-finetune paradigm as an effective post-training pipeline for TSFMs. To the best of our knowledge, this is the first work that explores pruning TSFMs to enhance forecasting performance. Remarkably, we find that up to 90% parameters can be removed while simultaneously improving performance.
Demonstrating performance and efficiency gains: we conduct extensive experiments on various TSFMs and datasets. The results demonstrate that our minimalist pruning approach frequently enables TSFMs to surpass state-of-the-art baselines in full-shot forecasting scenarios. Additionally, our structured pruning technique delivers up to 7 speedup in inference efficiency.
2 Related Works
2.1 Time Series Foundation Models
Recently, increasing efforts have been devoted to pre-training TSFMs on large-scale time series data for a versatile starting point of various prediction tasks. Timer [21] and Timer-XL [20] [2] stack multiple Transformer layers with causal attention to capture temporal correlations. Furthermore, several specialization modules are designed to handle the heterogeneity of time series. For frequency-specific specialization, Moirai [28] differentiates input/output projection layers among frequencies, while TTM [7] and TimesFM [6] include categorical frequency embeddings as input. Time-MoE [26] and Moirai-MoE [19] adopt a sparse mixture-of-experts (MoE) design where tokens are processed by different expert feed-forward networks (FFNs). Though only a few experts are activated for each token, all experts can be used and fine-tuned on the whole dataset.
While some TSFMs claim superior zero-shot or full-shot performance over specialized models like PatchTST [23], their comparisons set the lookback length to 96 for specialized models while thousands for TSFMs. Li et al. demonstrate that specialized models with well-tuned hyperparameters outperform TSFMs, while they did not tune TSFM hyperparameters. In this work, after extensively tuning hyperparameters of all models based on the validation set, our benchmark results confirm that TSFMs, though equipped with specialization modules, still need fine-tuning for higher downstream performance. Also, we find that applying parameter-efficient fine-tuning [13] to TSFMs delivers similar performance improvement compared with full fine-tuning (see Appendix C.2).
2.2 Model Pruning
Pruning [5], a conventional technique for network compression and acceleration, is increasingly applied to Large Language Models (LLMs) for efficient inference. Based on pruning units, pruning has two categories. Unstructured pruning removes unimportant elements of weight matrices, while the sparse matrices require specialized hardware or software optimizations for acceleration [34]. Structured pruning removes entire components (e.g., channels and attention heads with the lowest importance) and is hardware-friendly. Numerous pruning metrics have been proposed to determine whether a pruning unit is essential and should be pruned or not. Various metrics guide pruning; for example, LLM-Pruner [22] assesses a component’s importance by its impact on loss. Typically, LLMs after structure pruning struggle to regain original performance even after retraining [22, 3, 34]. The potential reason behind lossy compression of LLMs is that various neurons can preserve language skills and vast factual knowledge, which is critical to commonsense questions.
Crucially, the motivation of our work differs from the predominant use of pruning for LLMs. Our main purpose is not to pursue inference efficiency. Instead, we are devoted to improving the fine-tuning performance of TSFMs by pruning neurons that are less relevant to the downstream task. The rationale is that time series are characterized by strong heterogeneity across domains and frequencies [19], and a specific task may only need to activate a certain fraction of model parameters [26]. Moreover, due to the nature of non-stationarity, time series usually exhibit various data patterns over time, and there are scarce latent features that maintain long-term effectiveness for time series forecasting. Thus, pruning has the potential to focus TSFMs on the most critical channels for a given task.
3 Preliminary
In this section, we first formalize the notations of key modules in Transformers, which constitute the backbone of most existing TSFMs. Then, we conduct a preliminary study on TSFMs in aspects of attention head outputs and FFN activations. Additionally, analysis of parameter magnitudes is provided in Appendix B.1.
3.1 Transformer Architecture
Multi-Head Attention (MHA).
Typically, TSFMs organize time series observations as a sequence of tokens (a.k.a., patches). Let denote the number of input tokens and denote the token embeddings. Suppose there are attention heads and each head has dimensions. For each , let denote key, query, value projections for the -th head, and for output projections. Formally, MHA computation can be defined as below:
where is the output of the -th head.
Feed-Forward Network (FFN).
Each FFN involves at least two linear layers parameterized by . FFN computation can be written as
where is the activation function (e.g., ReLU, GELU, SwiGLU) that can deactivate some intermediate hidden states as zeros or tiny negative values. TTM [7], though not based on Transformers, also incorporates MLPs with GELU activation. For the -th intermediate FFN channel, the probability of producing positive activations over a dataset can reflect its relevance to the downstream task.
Residual Connection.
To enable deeper architectures and stabilize training, Transformer employs residual connections and normalizations. Formally, the output of a Transformer can be written as
where can be implemented by layer normalization or RMSNorm.
3.2 Empirical Analysis of TSFMs
According to the theory of bias-variance trade-offs [9], a large model can induce high variance risks in predictions, which seems contradictory to the skilled zero-shot performance of TSFMs on unseen tasks. To ascertain the mechanism of TSFMs, we conduct an investigation into TSFMs through the lens of attention heads and activations in FFNs. We perform 96-step forecasting on the Weather [29] and ETTm1 [33] datasets by PatchTST and seven pre-trained TSFMs, including TTM [7], Time-MoE, Time-MoE [26], Moirai, Moirai [28], Chronos-bolt [2], and TimesFM-2.0-500M [6]. The key statistics of the TSFMs are provided in Table 1.
| Model | TTMA | Time-MoE | Time-MoE | Moirai | Moirai | Chronos | TimesFM | PatchTST |
| Architecture | MLP | Decoder- Only | Decoder- Only | Encoder- Only | Encoder- Only | Encoder- Decoder | Decoder- Only | Encoder- Only |
| Model Size | 5M | 113M | 453M | 91M | 311M | 205M | 500M | 500K |
| #Layer | 6 | 12 | 12 | 12 | 24 | 24 | 50 | 3 |
| #Head | \ | 12 | 12 | 12 | 16 | 12 | 16 | 12 |
| 26384 | 384 | 768 | 768 | 1024 | 1024 | 1280 | 128 | |
| 52768 | 1536 | 3072 | 3072 | 4096 | 3072 | 1280 | 256 |
3.2.1 Insignificant Outputs of Attention Heads
Empirically, we find that there are a few attention heads in TSFMs that usually produce insignificant modifications on residuals. For each head, we calculate its relative output norm averaged over the downstream dataset, as defined below.
Definition 1
Given the -th head output and the residual of each input token, the average relative output norm of the -th head is defined as over a downstream dataset, which can reflect the importance of the -th head when modifying the residual.
Fig. 2 illustrates diverse distributions of average relative output norms. Compared to PatchTST where all heads actively contribute relative norm, a few heads of pre-trained TSFMs give nearly zero norms, and larger TSFM variants tend to exhibit greater sparsity. For instance, 50% (resp. 30%) heads of Time-MoE (resp. Moirai) produce outputs with negligible relative norms (), while their smaller counterparts have fewer insignificant heads. Nevertheless, with parameters more than Time-MoE and Moirai, Chronos exhibits lower sparsity, suggesting that the sparsity degree varies with both model sizes and architectures.
[skip=0pt]
3.2.2 Sparse Activations in FFNs
Shi et al. and Liu et al. found that MoE-based TSFMs tend to allocate different expert FFNs for samples from diverse datasets, and sparsely activated experts can address the heterogeneity of time series. We assume that there is also a sparse activation pattern at the granularity of FFN channels.
Definition 2
The activation probability of the -th intermediate channel of FFNs is defined as the frequency of its activations (denoted by ) being positive, i.e., , over the dataset.
Fig. 3 highlights distinct FFN activation patterns. PatchTST utilizes their FFN channels relatively densely, with all channels active at least 20% computations. By contrast, several TSFMs exhibit extreme sparsity: 20-60% of FFN channels in Chronos and Time-MoE models remain entirely inactive on the downstream data. Notably, even non-MoE architectures like Chronos and Moirai demonstrate this ability to selectively activate FFN parameter subgroups for specific tasks. This sparsity tends to increase with model size, as seen with Moirai being sparser than Moirai. Even the smaller TTM (5M parameters) shows some FFN channels with low activation probabilities. Defining channels with <5% activation probability as unimportant, Fig. 4 further illustrates that significant FFN sparsity patterns are prevalent across different layers within TSFMs.
According to our analysis on attention heads and FFN activations, most pre-trained TSFMs exhibit inherent sparsity during inference, with a substantial portion of parameters being either redundant or inactive. Moreover, TSFMs exhibit different sparsity patterns over the Weather and ETTm1 datasets. We hypothesize that their impressive zero-shot forecasting performance stems from the ability to selectively activate a task-relevant subset of parameters, leveraging prior knowledge. This encourages us to explore more fine-grained sparsity estimations and universal pruning methods.
4 Methodology
In this section, we first define the pruning unit and its importance score, followed by a progressive pruning and fine-tuning method to specialize TSFMs for downstream tasks.
4.1 Pruning Unit
For a comprehensive removal of task-irrelevant parameters, we define a pruning unit (i.e., the minimal pruning elements) as the input/output channel of any linear layer, covering self-attention modules and FFNs. Formally, a linear layer is parameterized by a transformation matrix and (optionally) a bias term , where and are the dimension of the input and output channels, respectively. Additionally, we apply two binary mask vectors and to the input and output of , respectively, which are initialized as . Given an input , the masked output is defined as below:
where corresponds to the element-wise product. We can prune the -th row of by setting , the -th element of , to zero. Likewise, the -th column of and the -th element of will be pruned if . It is worth mentioning that due to the residual connection and multi-head attention in Transformers, masking the output channels of a given layer is not equivalent to masking the input channels in the subsequent layer. Therefore, we treat the masking operations on input and output channels independently.
4.2 Importance Score of Pruning Unit
From the perspective of forecasting loss, a redundant channel, if pruned, would result in negligible loss changes. Following prior works [22], we define the channel-wise importance score as follows:
Definition 3
The importance score of channel , denoted as , is defined by , where denote a one-hot vector with the -th element equal to .
For computational efficiency, we derive an approximation of by:
where is the number of samples involved in calculating forecast loss , and is the forecasting loss of the -th sample. In Eq. (5), we apply the second-order Taylor expansion to the forecasting loss to efficiently evaluate the loss change. In Eq. (6), we employ Fisher information to approximate the second-order derivative by the square of the first-order derivative [18].
The derivative w.r.t. is computed by In a special case when is always a zero activation, the -th input channel has a zero importance score and will be pruned. The derivative of can be computed by The output channels of an attention head will also have low importance scores if the outputs are marginal and the loss derivatives are small. On top of the inherent sparsity in TSFMs, the importance score metric can also measure other redundant channels whose outputs are of great value but get counteracted by other channels or activation functions in deeper layers. Therefore, we adopt the importance score for a comprehensive pruning to extract the critical network structure.
Analysis.
Fig. 5 illustrates the distribution of channel importance scores over the Weather dataset. The long-tailed distribution of these scores within TSFMs demonstrates significant redundancy in channels. This channel redundancy is notably more pronounced in larger TSFM variants. This TSFM behavior caters to the nature of time series forecasting, where the presence of concept drift [10, 14, 30, 32, 31] implies that enduringly effective data patterns are scarce. Consequently, structured pruning emerges as a compelling strategy to selectively restrain the number of exploited latent features.
4.3 Progressive Pruning and Then Finetune
To avoid over-pruning at one time, it is desirable to remove a small number of channels at one pruning step and re-calculate the loss at the next step. Given the high expense of inferring the whole training set repeatedly, we adopt pruning based on batches. First, sampling the -th batch, we estimate the current importance score by Eq. (6). Second, given previously calculated scores, we estimate an exponential moving average (EMA) by , where is a hyperparameter. Third, globally across all model layers, we rank in ascending order and remove unpruned channels with the lowest importance based on a predefined pruning ratio. Finally, we obtain a pruned model after one or a few epochs over the training set. Optionally for acceleration, one can first prune out insignificant attention heads and sparse FFN channels discovered in the forward pass based on a predefined threshold. Then, we involve gradients to compute importance scores of other channels for a fine-grained pruning. After pruning, we fine-tune the remaining parameters on the downstream data.
5 Experiments
In this section, we present extensive experiments to evaluate the effectiveness of structured pruning when adapting TSFMs to downstream tasks. We compare full-shot forecasting performance with and without pruning, analyze the resulting sparsity patterns and inference speedups, and assess the transferability of the pruned TSFMs to other datasets. Additional experimental results are provided in Appendix C, including ablation studies on the tuning methods (Table 7) and the pruning methods (Table 8).
5.1 Experimental Settings
Foundation Models.
Our experiments are conducted on a range of pre-trained Transformer-based time series forecasting models (TSFMs), including Time-MoE111https://huggingface.co/Maple728/TimeMoE-50M, Time-MoE222https://huggingface.co/Maple728/TimeMoE-200M, Timer-XL333https://huggingface.co/thuml/timer-base-84m, Moirai444https://huggingface.co/Salesforce/Moirai-1.0-R-base, Moirai555https://huggingface.co/Salesforce/Moirai-1.0-R-large, Chronos-bolt666https://huggingface.co/amazon/chronos-bolt-base, and TTM777https://huggingface.co/ibm-research/ttm-research-r2. We inherit the open-source codes and pre-trained checkpoints from their official repositories.
Datasets.
We evaluate long-term forecasting performance on three widely used benchmarks: ETT [33], Weather [29], and Electricity [15]. Dataset statistics and train/validation/test splits are summarized in Table 5. Pruning and fine-tuning are performed on the training sets, and we report forecasting accuracy on the test sets using Mean Squared Error (MSE) and Mean Absolute Error (MAE) over horizons of 96, 192, 336, and 720 steps.
Implementation Details.
To ensure fair comparison, we carefully tune all models’ hyperparameters (e.g., learning rate, pruning ratio) on the validation set. For autoregressive TSFMs, a single model is trained and pruned per dataset and reused across all forecasting lengths. In contrast, for non-autoregressive models (e.g., Moirai and TTM), we train separate models optimized for each forecasting horizon. Further details on the hyperparameter search process are provided in Appendix A.2 to facilitate reproducibility. For simplicity, our main pruning strategy relies solely on importance scores. Additional results using alternative pruning criteria are presented in the supplementary material.
5.2 Forecasting Performance
As shown in Table 2, our prune-then-finetune paradigm achieves higher predictive performance than directly fine-tuning TSFMs in 83% of forecasting tasks. The prediction errors are reduced by an average of 4.4%, 4.0%, 2.5%, 3.2%, 1.3%, 3.3%, 1.0% for Time-MoE, Time-MoE, Timer-XL, Moirai, Moirai, Chronos, TTMA, respectively. The maximum reduction is 22.8% (for Time-MoE on ETTm2), while the maximum performance drop is capped at only 1%. Notably, our method significantly increases the win rate of several TSFMs against PatchTST, with the strong baseline surpassed on all benchmarks, showcasing the remarkable potential of structured pruning. As for autoregressive predictions of TimeMoE, Timer-XL, and Chronos, their performance gains from pruning become more pronounced with longer forecasting horizons, due to reduced error accumulation. We posit that our pruning method goes beyond the inherent sparsity of TSFMs by deactivating more channels to regularize exploiting latent features, while fine-tuning the original model may mistakenly focus on outdated or noisy features. Different from random dropout, pruning focuses on a task-relevant subspace of parameters consistently across fine-tuning and inference.
Model TimeMoE TimeMoE Timer-XL Moirai Moirai Chronos TTMA PatchTST Method FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT Tasks MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE 96 0.340 0.377 0.340 0.376 0.331 0.372 0.329 0.372 0.365 0.393 0.353 0.391 0.367 0.396 0.366 0.396 0.371 0.391 0.370 0.391 0.366 0.374 0.366 0.375 0.359 0.395 0.358 0.393 0.376 0.396 192 0.380 0.408 0.376 0.404 0.371 0.398 0.369 0.399 0.400 0.416 0.389 0.413 0.405 0.418 0.404 0.418 0.416 0.416 0.413 0.416 0.418 0.405 0.417 0.405 0.392 0.417 0.391 0.410 0.399 0.416 336 0.406 0.432 0.399 0.425 0.392 0.413 0.389 0.411 0.419 0.431 0.409 0.426 0.435 0.434 0.429 0.430 0.455 0.431 0.455 0.432 0.453 0.424 0.449 0.422 0.407 0.430 0.402 0.424 0.418 0.432 720 0.437 0.469 0.424 0.457 0.403 0.438 0.387 0.419 0.454 0.465 0.433 0.455 0.465 0.459 0.427 0.441 0.475 0.457 0.454 0.444 0.452 0.441 0.451 0.437 0.448 0.467 0.449 0.465 0.450 0.469 Avg 0.391 0.422 0.385 0.416 0.374 0.405 0.369 0.400 0.410 0.426 0.396 0.421 0.418 0.427 0.407 0.421 0.429 0.424 0.423 0.421 0.422 0.411 0.421 0.410 0.402 0.427 0.400 0.423 0.411 0.428 ETTh1 #p 63% 63% 67% 58%69% 36%39% 55% 91%93% 96 0.267 0.330 0.262 0.327 0.272 0.332 0.272 0.332 0.287 0.345 0.275 0.337 0.273 0.338 0.272 0.338 0.288 0.341 0.287 0.341 0.266 0.311 0.266 0.311 0.264 0.333 0.262 0.328 0.277 0.339 192 0.341 0.380 0.331 0.374 0.356 0.383 0.356 0.382 0.344 0.385 0.334 0.378 0.331 0.385 0.331 0.385 0.349 0.386 0.350 0.387 0.334 0.357 0.331 0.356 0.321 0.372 0.324 0.370 0.345 0.381 336 0.386 0.415 0.372 0.403 0.420 0.424 0.411 0.418 0.375 0.410 0.368 0.404 0.358 0.410 0.358 0.409 0.373 0.409 0.371 0.408 0.378 0.391 0.370 0.388 0.351 0.397 0.351 0.396 0.368 0.404 720 0.419 0.454 0.377 0.423 0.534 0.494 0.459 0.454 0.411 0.445 0.398 0.439 0.392 0.444 0.388 0.442 0.402 0.439 0.394 0.435 0.395 0.418 0.389 0.415 0.395 0.439 0.393 0.437 0.397 0.432 Avg 0.353 0.395 0.336 0.382 0.396 0.408 0.375 0.397 0.354 0.396 0.344 0.390 0.339 0.394 0.337 0.394 0.353 0.394 0.351 0.393 0.343 0.369 0.339 0.368 0.333 0.385 0.333 0.383 0.347 0.389 ETTh2 #p 82% 77% 68% 55%78% 27%36% 56% 93%98% 96 0.220 0.302 0.220 0.302 0.214 0.305 0.211 0.299 0.305 0.355 0.297 0.354 0.309 0.343 0.304 0.342 0.311 0.345 0.312 0.346 0.285 0.315 0.283 0.315 0.295 0.346 0.291 0.340 0.290 0.343 192 0.284 0.347 0.282 0.344 0.268 0.345 0.257 0.333 0.348 0.383 0.337 0.381 0.349 0.367 0.341 0.366 0.343 0.370 0.342 0.370 0.329 0.345 0.326 0.344 0.332 0.366 0.327 0.363 0.329 0.368 336 0.349 0.392 0.335 0.381 0.326 0.386 0.300 0.364 0.382 0.406 0.367 0.403 0.379 0.385 0.375 0.384 0.364 0.385 0.359 0.385 0.361 0.369 0.356 0.367 0.356 0.383 0.356 0.381 0.360 0.390 720 0.462 0.466 0.415 0.437 0.428 0.449 0.380 0.415 0.428 0.439 0.409 0.434 0.410 0.409 0.397 0.409 0.401 0.409 0.400 0.409 0.434 0.413 0.423 0.407 0.409 0.420 0.395 0.406 0.416 0.422 Avg 0.329 0.377 0.313 0.366 0.309 0.371 0.287 0.353 0.366 0.396 0.353 0.393 0.362 0.376 0.354 0.375 0.355 0.377 0.353 0.378 0.352 0.361 0.347 0.358 0.348 0.379 0.342 0.373 0.349 0.381 ETTm1 #p 92% 64% 81% 51%60% 51%52% 55% 86%99% 96 0.161 0.253 0.158 0.251 0.166 0.258 0.164 0.258 0.191 0.280 0.181 0.273 0.174 0.261 0.172 0.259 0.165 0.251 0.165 0.251 0.160 0.231 0.157 0.231 0.160 0.251 0.160 0.248 0.165 0.254 192 0.229 0.308 0.217 0.299 0.248 0.318 0.236 0.313 0.245 0.320 0.233 0.310 0.234 0.304 0.231 0.300 0.234 0.304 0.232 0.301 0.225 0.276 0.219 0.274 0.213 0.289 0.213 0.286 0.221 0.292 336 0.313 0.367 0.273 0.342 0.347 0.381 0.314 0.367 0.297 0.356 0.283 0.344 0.284 0.340 0.282 0.337 0.288 0.343 0.287 0.340 0.286 0.317 0.276 0.312 0.269 0.326 0.268 0.322 0.275 0.325 720 0.486 0.464 0.375 0.407 0.552 0.488 0.471 0.453 0.389 0.416 0.367 0.396 0.353 0.386 0.350 0.384 0.366 0.397 0.365 0.396 0.381 0.377 0.363 0.369 0.346 0.379 0.338 0.369 0.360 0.380 Avg 0.297 0.348 0.256 0.325 0.328 0.361 0.296 0.348 0.281 0.343 0.266 0.331 0.261 0.323 0.259 0.320 0.263 0.324 0.262 0.322 0.263 0.300 0.254 0.297 0.247 0.311 0.245 0.306 0.255 0.313 ETTm2 #p 75% 57% 64% 51%56% 35%38% 55% 86%99% 96 0.141 0.194 0.141 0.196 0.140 0.193 0.139 0.190 0.169 0.226 0.165 0.219 0.178 0.228 0.159 0.222 0.147 0.199 0.147 0.198 0.174 0.209 0.141 0.175 0.150 0.202 0.145 0.195 0.149 0.196 192 0.196 0.254 0.191 0.250 0.191 0.245 0.189 0.241 0.217 0.271 0.209 0.261 0.220 0.269 0.216 0.265 0.211 0.263 0.198 0.255 0.231 0.258 0.186 0.220 0.192 0.243 0.188 0.240 0.193 0.240 336 0.260 0.309 0.249 0.301 0.255 0.297 0.253 0.292 0.267 0.308 0.255 0.296 0.308 0.330 0.238 0.287 0.277 0.315 0.279 0.317 0.289 0.298 0.239 0.262 0.237 0.279 0.236 0.277 0.244 0.281 720 0.346 0.378 0.331 0.370 0.371 0.383 0.361 0.370 0.336 0.358 0.322 0.345 0.390 0.388 0.333 0.361 0.374 0.383 0.345 0.365 0.376 0.347 0.319 0.315 0.314 0.335 0.313 0.335 0.314 0.332 Avg 0.236 0.284 0.228 0.279 0.239 0.280 0.236 0.273 0.247 0.291 0.238 0.280 0.274 0.304 0.237 0.284 0.252 0.290 0.242 0.284 0.268 0.278 0.221 0.243 0.223 0.265 0.221 0.262 0.225 0.262 Weather #p 68% 88% 75% 20%62% 20%44% 55% 91%94% 96 - - - - - - - - 0.139 0.237 0.138 0.237 0.152 0.247 0.135 0.236 0.148 0.243 0.138 0.236 0.121 0.208 0.121 0.208 0.129 0.226 0.129 0.226 0.133 0.233 192 - - - - - - - - 0.158 0.256 0.157 0.255 0.170 0.263 0.162 0.263 0.172 0.265 0.164 0.265 0.139 0.226 0.139 0.225 0.147 0.243 0.148 0.244 0.150 0.248 336 - - - - - - - - 0.177 0.275 0.176 0.275 0.192 0.285 0.176 0.275 0.187 0.280 0.190 0.282 0.157 0.244 0.156 0.243 0.167 0.266 0.166 0.265 0.168 0.267 720 - - - - - - - - 0.219 0.313 0.217 0.312 0.220 0.307 0.210 0.305 0.219 0.309 0.215 0.310 0.199 0.283 0.195 0.278 0.202 0.297 0.199 0.295 0.202 0.295 Avg - - - - - - - - 0.173 0.270 0.172 0.270 0.184 0.276 0.171 0.270 0.182 0.274 0.177 0.273 0.154 0.240 0.153 0.239 0.161 0.258 0.161 0.258 0.163 0.261 Electricity #p 56% 40%42% 3%60% 53% 94%98% Win Rate 46.0% 68.0% 46.0% 52.0% 11.7% 26.7% 25.0% 33.3% 18.3% 20.0% 80.0% 95.0% 85.0% 93.3% • * “#p” is the proportion of the remaining parameters after pruning relative to the original ones. • * “Win Rate” is based on comparisons between each TSFM and PatchTST over all datasets. • * Electricity included in the pre-training of Time-MoE is not evaluated, denoted by a dash (–). • * We present performance comparison about TimesFM in Table 8.
5.3 Sparsification Analysis
As shown in Fig. 6, we compare the sparsification results in terms of the pruning ratios across input and output channels. Although both models consist of the same number of Transformer layers, Moirai adopts an encoder-only architecture, while Chronos follows an encoder-decoder design. Moirai generally exhibits more aggressive pruning in its deeper layers, with non-negligible pruning also observed in its shallow layers. In contrast, pruning in Chronos is predominantly concentrated in the decoder (i.e., the last 12 layers), while the encoder remains largely intact. We attribute this to its use of a much smaller patch size for time series tokenization, which requires the encoder to capture and aggregate more contextual information across tokens.
It is also important to note that channel pruning occurs across various layer types. Importance-score-based pruning successfully removes certain attention heads and sparsely activated FFN channels, as evidenced by the pruned input channels of and , respectively. Furthermore, additional sparsification is introduced in other projection matrices, including , , , and . The observed improvements in fine-tuning performance confirm the effectiveness of our method in systematically identifying and pruning redundant components using importance-based criteria.
[figure]skip=-1pt
5.4 Inference Efficiency
As shown in Table 3, in addition to the improvement in predictive performance, our proposed prune-then-finetune approach also gives a bonus in the speed of inference. We evaluate the computational efficiency of pruned TSFMs using the Weather benchmark. At each time step, all variates are processed as a batch, and we report the average CPU runtime during forecasting 720 steps, measured over 50 batches. Our results show that the pruned TSFMs achieve up to a 7.4× speedup, with the largest gain observed for Moirai.
| Original | Pruned | Speedup | |
| TimeMoE | 72 | 70 | 1.0 |
| TimeMoE | 110 | 103 | 1.1 |
| Timer-XL | 0.18 | 0.16 | 1.1 |
| Moirai | 0.37 | 0.14 | 2.6 |
| Moirai | 1.86 | 0.25 | 7.4 |
| Chronos | 8.51 | 7.19 | 1.2 |
| TTMA | 0.01 | 0.01 | 1.0 |
5.5 Transferability
We also observe that the pruned model exhibits strong transferability across related downstream tasks. We study Chronos using the ETTh and ETTm datasets from the electricity domain, with results summarized in Table 4. Notably, with appropriate source–target task pairs, the pruned model transferred from the source domain can even outperform the original (unpruned) model on the target task. This suggests that sparsity patterns learned in one domain may generalize well to semantically related tasks. Consequently, in data-scarce settings, one can leverage similar datasets to guide pruning and architecture specialization, potentially leading to improved performance.
| Model | Chronos | |||
| Methods | Zero-shot | Prune | ||
| SourceTarget | MSE | MAE | MSE | MAE |
| ETTh1ETTh2 | 0.420 | 0.421 | 0.389 | 0.410 |
| ETTh2ETTh1 | 0.473 | 0.447 | 0.465 | 0.444 |
| ETTm1ETTm2 | 0.420 | 0.395 | 0.396 | 0.385 |
| ETTm2ETTm1 | 0.504 | 0.430 | 0.501 | 0.427 |
| ETTm1ETTh2 | 0.420 | 0.421 | 0.408 | 0.418 |
| ETTm2ETTh1 | 0.473 | 0.447 | 0.469 | 0.446 |
| ETTh1ETTm2 | 0.420 | 0.395 | 0.376 | 0.376 |
| ETTh2ETTm1 | 0.504 | 0.430 | 0.491 | 0.426 |
6 Conclusion
This paper sheds light on the critical challenge of adapting pre-trained TSFMs for superior performance on full-shot downstream forecasting tasks. Investigating inherent sparsity and redundancy within TSFMs, we suggest that the activated task-relevant subnetwork is valuable prior knowledge of a starting point for fine-tuning. We proposed a “prune-then-finetune” paradigm, employing structured pruning to identify and extract the critical subnetwork. Extensive experiments across seven TSFMs and six popular benchmarks compellingly demonstrated that our method significantly improves forecasting performance, enabling pruned TSFMs to surpass strong specialized baselines. This work pioneers structured pruning primarily for performance enhancement via architectural specialization, offering a promising pathway to unlock the full potential of TSFMs in practical applications.
Appendix A Implementation Details
A.1 Benchmark Details
| Dataset | Variate | Dataset Split | Frequency | Variate Types |
| ETTh1 | 7 | (8545, 2881, 2881) | Hourly | electricity load, oil temperature |
| ETTh2 | 7 | (8545, 2881, 2881) | Hourly | electricity load, oil temperature |
| ETTm1 | 7 | (34465, 11521, 11521) | 15 min | electricity load, oil temperature |
| ETTm2 | 7 | (34465, 11521, 11521) | 15 min | electricity load, oil temperature |
| Weather | 21 | (36792, 5271, 10540) | 10 min | temperature, wind speed, pressure, rain, etc. |
| Electricity | 321 | (12280, 1755, 3509) | Hourly | electricity load |
We evaluate the long-term forecasting performance across six well-established datasets, including the Weather [29], Electricity [15], and ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) [33]. A detailed description of each dataset is provided in Table 5. We use the full test set without the “drop-last” strategy used in PatchTST [23]. We borrow benchmark results of PatchTST from TFB [24] that carefully tuned the hyperparameters.
A.2 Experimental Details
We conduct extensive experiments based on 40 Nvidia V100 GPUs, 32 Nvidia 3090 GPUs, and Intel(R) Xeon(R) Platinum 8163 CPUs. The random seed is set to 1. The search grid of hyperparameters is provided in Table 6. Particularly, we set the fine-tuning epoch to 1 for Time-MoE, Chronos, and Moirai; otherwise, these models after multi-epoch fine-tuning usually achieve lower validation MSE but higher test MSE due to overfitting. To reduce tuning cost, we do not run all combinations of hyperparameters. For instance, we tune the hyperparameters of pruning and fine-tuning separately. As for pruning, we first tune the coefficient of exponential moving average that is optimal for a specific pruning ratio per epoch. Given four pruning ratios, we obtain four pruned models and tune the fine-tuning learning rates for each model. Alternatively, we can also select one pruned model that performs the best on the validation set and tune the learning rate of only a single model. The pruned models always follow the same batch size for fine-tuning the original model. The hyperparameter tuning for the most expensive model (e.g., Moirai) on the largest Electricity dataset consumes about 1 day.
| Time-MoE | TTM | Time-XL | Moirai | Chronos | |
| Learning rate | {1e-2, 1e-3, 1e-4, 1e-5, 1e-6, 1e-7, 1e-8, 1e-9} | ||||
| Context length | 4096 | 1536 | 2880 | {2048, 3072, 4096} | 2048 |
| Patch size | 1 | 1 | 96 | {32, 64, 128} | 16 |
| Batch size for fine-tuning | {32, 64} | {256, 512} | |||
| Max. fine-tuning epoch | 1 | 10 | 10 | 1 | 1 |
| Patience for early stop | 1 | ||||
| Pruning ratio per epoch | {1%, 2%, 5%, 10%, 15%, 20%} | ||||
| Exponential moving average | {0.1, 0.2, 0.4, 0.5, 0.6, 0.8} | ||||
| Batch size for pruning | 8192 | ||||
| Max. pruning epoch | 10 | 1 | 10 | 10 | 1 |
Appendix B Additional Empirical Study
B.1 Insignificant Weight Magnitudes
On top of the analysis of sparse attention heads and FFN channels, another question is whether all parameters of TSFMs are of vital importance. To ascertain it, we analyze the distribution of parameter magnitudes, which are defined as the absolute value of each element in weight matrices. For a comparable study across models, we normalize the magnitude by dividing each by the maximum value, ensuring all values fall within the [0, 1] range. As shown in Fig. 7(a), Chronos and Moirai, based on Transformer encoders, have a considerable proportion of tiny parameter weights. In contrast, PatchTST has a nearly uniform distribution of weight magnitudes, though it is encoder-based as well. As for decoder-only TSFMs, Timer-XL and Time-MoE also have a skewed distribution of weight magnitudes. As for the vector-wise magnitude of parameter weights (i.e., the rows or columns of weight matrices), we depict the cumulative distribution in Fig. 7(b), where the difference between TSFMs and PatchTST becomes more significant. These findings suggest that zero-shot forecasting may be based on a portion of parameters in TSFMs instead of all parameters. Nevertheless, it is worth mentioning that pruning based on magnitude is task-agnostic and may result in suboptimal downstream performance. This is why we adopt loss change to compute importance scores.
B.2 Importance Scores over the ETTm1 dataset
We depict the distribution of importance scores over the ETTm1 dataset in Fig. 8. Compared with results on Weather as shown in Fig. 5, the redundancy patterns of the pre-trained TSFMs vary on the datasets, indicating that there are some task-specific parameters.
Appendix C Additional Experimental Results
C.1 Sparsification Visualization
We visualize more sparsification results in Fig. 10. Moirai and Timer-XL exhibit sparsification in input channels of output projections and intermediate channels of FFNs, which confirms that pruning based solely on importance scores can handle the inherent sparsity of TSFMs. In contrast, the sparsification of Time-MoE is distinct by mainly pruning the FFNs. This can be attributed to the fact that the usage of expert FFNs in MoE models is less frequent than the multi-head attention modules, resulting in less contribution to the forecast loss over a batch. Moreover, the multiple experts may share redundancy with each other, and pruning an expert may not lead to significant loss changes. In our supplementary materials, we will provide results of pruning the attention heads based on a threshold of the relative output norm.
C.2 Performance of Other Tuning Methods
| Models | Chronos | TTMA | |||||||||||||||
| Methods | LoRA | Full FT | Prune+LoRA | Prune+Full FT | LoRA | Full FT | Prune+LoRA | Prune+Full FT | |||||||||
| Datasets | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
| 96 | 0.366 | 0.374 | 0.366 | 0.374 | 0.366 | 0.375 | 0.366 | 0.375 | 0.360 | 0.396 | 0.359 | 0.395 | 0.358 | 0.393 | 0.358 | 0.393 | |
| 192 | 0.420 | 0.405 | 0.418 | 0.405 | 0.417 | 0.405 | 0.417 | 0.405 | 0.394 | 0.420 | 0.392 | 0.417 | 0.388 | 0.411 | 0.391 | 0.410 | |
| 336 | 0.456 | 0.425 | 0.453 | 0.424 | 0.450 | 0.422 | 0.449 | 0.422 | 0.409 | 0.433 | 0.407 | 0.430 | 0.406 | 0.429 | 0.402 | 0.424 | |
| 720 | 0.454 | 0.441 | 0.452 | 0.441 | 0.451 | 0.437 | 0.451 | 0.437 | 0.448 | 0.471 | 0.448 | 0.467 | 0.447 | 0.464 | 0.449 | 0.465 | |
| ETTh1 | Avg. | 0.424 | 0.411 | 0.422 | 0.411 | 0.421 | 0.410 | 0.421 | 0.410 | 0.403 | 0.430 | 0.402 | 0.427 | 0.400 | 0.424 | 0.400 | 0.423 |
| 96 | 0.267 | 0.312 | 0.266 | 0.311 | 0.266 | 0.311 | 0.266 | 0.311 | 0.264 | 0.333 | 0.264 | 0.333 | 0.262 | 0.329 | 0.262 | 0.328 | |
| 192 | 0.335 | 0.358 | 0.334 | 0.357 | 0.333 | 0.356 | 0.331 | 0.356 | 0.321 | 0.372 | 0.321 | 0.372 | 0.324 | 0.371 | 0.324 | 0.370 | |
| 336 | 0.379 | 0.392 | 0.378 | 0.391 | 0.376 | 0.389 | 0.370 | 0.388 | 0.351 | 0.397 | 0.351 | 0.397 | 0.352 | 0.396 | 0.351 | 0.396 | |
| 720 | 0.395 | 0.418 | 0.395 | 0.418 | 0.392 | 0.414 | 0.389 | 0.415 | 0.395 | 0.439 | 0.395 | 0.439 | 0.389 | 0.432 | 0.393 | 0.437 | |
| ETTh2 | Avg. | 0.344 | 0.370 | 0.343 | 0.369 | 0.342 | 0.368 | 0.339 | 0.368 | 0.333 | 0.385 | 0.333 | 0.385 | 0.332 | 0.382 | 0.333 | 0.383 |
| 96 | 0.285 | 0.315 | 0.285 | 0.315 | 0.285 | 0.315 | 0.283 | 0.315 | 0.295 | 0.346 | 0.295 | 0.346 | 0.289 | 0.339 | 0.291 | 0.340 | |
| 192 | 0.330 | 0.344 | 0.329 | 0.345 | 0.329 | 0.344 | 0.326 | 0.344 | 0.334 | 0.368 | 0.332 | 0.366 | 0.327 | 0.365 | 0.327 | 0.363 | |
| 336 | 0.363 | 0.368 | 0.361 | 0.369 | 0.362 | 0.367 | 0.356 | 0.367 | 0.358 | 0.384 | 0.356 | 0.383 | 0.357 | 0.382 | 0.356 | 0.381 | |
| 720 | 0.434 | 0.409 | 0.434 | 0.413 | 0.430 | 0.405 | 0.423 | 0.407 | 0.394 | 0.409 | 0.409 | 0.420 | 0.396 | 0.405 | 0.395 | 0.406 | |
| ETTm1 | Avg. | 0.353 | 0.359 | 0.352 | 0.361 | 0.352 | 0.358 | 0.347 | 0.358 | 0.345 | 0.377 | 0.348 | 0.379 | 0.342 | 0.373 | 0.342 | 0.373 |
| 96 | 0.161 | 0.232 | 0.160 | 0.231 | 0.159 | 0.232 | 0.157 | 0.231 | 0.161 | 0.254 | 0.160 | 0.251 | 0.161 | 0.248 | 0.160 | 0.248 | |
| 192 | 0.227 | 0.277 | 0.225 | 0.276 | 0.223 | 0.276 | 0.219 | 0.274 | 0.213 | 0.290 | 0.213 | 0.289 | 0.212 | 0.286 | 0.213 | 0.286 | |
| 336 | 0.287 | 0.318 | 0.286 | 0.317 | 0.280 | 0.315 | 0.276 | 0.312 | 0.262 | 0.323 | 0.269 | 0.326 | 0.266 | 0.320 | 0.268 | 0.322 | |
| 720 | 0.383 | 0.379 | 0.381 | 0.377 | 0.368 | 0.372 | 0.363 | 0.369 | 0.341 | 0.373 | 0.346 | 0.379 | 0.357 | 0.379 | 0.338 | 0.369 | |
| ETTm2 | Avg. | 0.265 | 0.302 | 0.263 | 0.300 | 0.258 | 0.299 | 0.254 | 0.297 | 0.244 | 0.310 | 0.247 | 0.311 | 0.249 | 0.308 | 0.245 | 0.306 |
| 96 | 0.156 | 0.190 | 0.174 | 0.209 | 0.152 | 0.191 | 0.141 | 0.175 | 0.153 | 0.208 | 0.150 | 0.202 | 0.145 | 0.195 | 0.145 | 0.195 | |
| 192 | 0.198 | 0.233 | 0.231 | 0.258 | 0.203 | 0.239 | 0.186 | 0.220 | 0.192 | 0.245 | 0.192 | 0.243 | 0.189 | 0.241 | 0.188 | 0.240 | |
| 336 | 0.248 | 0.272 | 0.289 | 0.298 | 0.267 | 0.283 | 0.239 | 0.262 | 0.240 | 0.281 | 0.237 | 0.279 | 0.238 | 0.278 | 0.236 | 0.277 | |
| 720 | 0.324 | 0.323 | 0.376 | 0.347 | 0.330 | 0.329 | 0.319 | 0.315 | 0.301 | 0.323 | 0.314 | 0.335 | 0.314 | 0.338 | 0.313 | 0.335 | |
| Weather | Avg. | 0.232 | 0.255 | 0.268 | 0.278 | 0.238 | 0.261 | 0.221 | 0.243 | 0.222 | 0.264 | 0.223 | 0.265 | 0.222 | 0.263 | 0.221 | 0.262 |
Parameter-Efficient Fine-Tuning (PEFT) has become a popular approach to efficient LLM fine-tuning by reducing the number of trainable parameters. LoRA [13] is one of the most prevalent PEFT methods that learns two additional low-rank weight matrices as an adjustment to the original full-rank weight matrix at each layer. In this section, we apply LoRA to Chronos and TTMA with the rank hyperparameter chosen from {32, 64, 128}. As shown in Table 7, we do not observe considerable performance improvement of LoRA compared to full fine-tuning in most cases.
It is noteworthy that our prune-then-finetune paradigm is orthogonal to any model tuning methods. We conduct additional experiments by incorporating LoRA into our paradigm. Concretely, the dimensions of LoRA matrices will depend on the remaining input/output dimensions of each linear transformation matrix, instead of the original ones. Similar to the performance gap between full fine-tuning and LoRA over the original TSFMs, applying full fine-tuning to pruned TSFMs is usually better than applying LoRA to pruned TSFMs. As for comparisons between LoRA and Prune+LoRA, the pruned ones can yield better performance except for a few cases. It is noteworthy that most GPU memory overhead during time series forecasting is caused by self-attention computations over long contexts, rather than the model parameters and the optimizer. Since the model sizes of existing TSFMs are much smaller than LLMs, parameter-efficient fine-tuning methods do not significantly reduce memory. Thus, we would like to adopt full fine-tuning for better performance.
C.3 Performance of Variant Pruning Methods
In light of the inherent sparsity of TSFMs, we introduce a variant of TSFM pruning, denoted as Prune. This method prunes TSFMs based on statistics during the forward pass, specifically including the relative output norm of attention heads and the activation probability of FFN channels. We collect these statistics over the whole training set and subsequently remove insignificant heads and channels whose statistics fall below a predefined threshold. The threshold for relative output norm is selected from {0%, 0.5%, 1%, 2%}, while that for activation probabilities is selected from {0%, 1%, 2%, 5%}. As a counterpart, the pruning method based on importance scores is denoted as Prune.
Table 8 presents forecasting performance of TSFMs specialized by variant pruning methods and full fine-tuning. The method “Prune+FT” usually achieves better performance than fine-tuning, which confirms the effectiveness of architecture specialization. Nevertheless, “Prune+FT” cannot outperform “Prune+FT” in most cases, indicating that only pruning attention heads or intermediate FFN channels is not sufficient for specialization. In a few cases, “Prune+FT” underperforms “Prune+FT”, e.g., given Time-MoE on ETTh1 where “Prune+FT” removes several attention heads while “Prune+FT” primarily focus on FFNs. We conjecture that a more comprehensive pruning strategy, combining both statistical and importance score-based criteria, could potentially lead to better performance.
Models TimeMoEbase TimesFM Chronosbase TTMA Methods FT Prune+FT Prune+FT FT Prune+FT Prune+FT FT Prune+FT Prune+FT FT Prune+FT Prune+FT Data MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE 96 0.340 0.377 0.339 0.374 0.340 0.376 0.384 0.391 0.385 0.391 0.383 0.393 0.366 0.374 0.366 0.374 0.366 0.375 0.359 0.395 0.359 0.395 0.358 0.393 192 0.380 0.408 0.374 0.400 0.376 0.404 0.417 0.412 0.419 0.412 0.414 0.413 0.418 0.405 0.418 0.405 0.417 0.405 0.392 0.417 0.390 0.414 0.391 0.410 336 0.406 0.432 0.393 0.415 0.399 0.425 0.430 0.422 0.433 0.423 0.419 0.422 0.453 0.424 0.453 0.424 0.449 0.422 0.407 0.430 0.408 0.431 0.402 0.424 720 0.437 0.469 0.406 0.440 0.424 0.457 0.432 0.441 0.434 0.441 0.429 0.444 0.452 0.441 0.451 0.437 0.451 0.437 0.448 0.467 0.437 0.461 0.449 0.465 ETTh1 Avg. 0.391 0.422 0.378 0.407 0.385 0.416 0.416 0.417 0.418 0.417 0.411 0.418 0.422 0.411 0.422 0.410 0.421 0.410 0.402 0.427 0.399 0.425 0.400 0.423 96 0.267 0.330 0.283 0.345 0.262 0.327 0.268 0.324 0.267 0.323 0.268 0.325 0.266 0.311 0.264 0.311 0.266 0.311 0.264 0.333 0.263 0.332 0.262 0.328 192 0.341 0.380 0.366 0.403 0.331 0.374 0.329 0.367 0.327 0.364 0.330 0.368 0.334 0.357 0.329 0.355 0.331 0.356 0.321 0.372 0.321 0.372 0.324 0.370 336 0.386 0.415 0.429 0.448 0.372 0.403 0.361 0.391 0.359 0.389 0.360 0.392 0.378 0.391 0.367 0.387 0.370 0.388 0.351 0.397 0.350 0.397 0.351 0.396 720 0.419 0.454 0.559 0.530 0.377 0.423 0.384 0.413 0.379 0.410 0.383 0.414 0.395 0.418 0.384 0.413 0.389 0.415 0.395 0.439 0.395 0.439 0.393 0.437 ETTh2 Avg. 0.353 0.395 0.409 0.432 0.336 0.382 0.336 0.374 0.333 0.372 0.335 0.375 0.343 0.369 0.336 0.367 0.339 0.368 0.333 0.385 0.332 0.385 0.333 0.383 96 0.220 0.302 0.221 0.301 0.220 0.302 0.321 0.338 0.316 0.336 0.298 0.338 0.285 0.315 0.285 0.315 0.283 0.315 0.295 0.346 0.298 0.347 0.291 0.340 192 0.284 0.347 0.278 0.343 0.282 0.344 0.366 0.365 0.359 0.363 0.340 0.365 0.329 0.345 0.329 0.344 0.326 0.344 0.332 0.366 0.330 0.366 0.327 0.363 336 0.349 0.392 0.333 0.382 0.335 0.381 0.390 0.385 0.382 0.382 0.368 0.385 0.361 0.369 0.361 0.368 0.356 0.367 0.356 0.383 0.358 0.384 0.356 0.381 720 0.462 0.466 0.427 0.444 0.415 0.437 0.434 0.415 0.423 0.411 0.419 0.420 0.434 0.413 0.430 0.409 0.423 0.407 0.409 0.420 0.392 0.405 0.395 0.406 ETTm1 Avg. 0.329 0.377 0.315 0.368 0.313 0.366 0.378 0.376 0.370 0.373 0.356 0.377 0.352 0.361 0.351 0.359 0.347 0.358 0.348 0.379 0.345 0.376 0.342 0.373 96 0.161 0.253 0.161 0.250 0.158 0.251 0.169 0.250 0.168 0.249 0.166 0.251 0.160 0.231 0.160 0.233 0.157 0.231 0.160 0.251 0.159 0.251 0.160 0.248 192 0.229 0.308 0.226 0.299 0.217 0.299 0.232 0.295 0.229 0.293 0.225 0.294 0.225 0.276 0.226 0.278 0.219 0.274 0.213 0.289 0.215 0.294 0.213 0.286 336 0.313 0.367 0.305 0.353 0.273 0.342 0.291 0.335 0.287 0.334 0.279 0.332 0.286 0.317 0.287 0.318 0.276 0.312 0.269 0.326 0.271 0.332 0.268 0.322 720 0.486 0.464 0.471 0.449 0.375 0.407 0.367 0.389 0.368 0.390 0.358 0.387 0.381 0.377 0.381 0.377 0.363 0.369 0.346 0.379 0.340 0.379 0.338 0.369 ETTm2 Avg. 0.297 0.348 0.291 0.338 0.256 0.325 0.265 0.317 0.263 0.317 0.257 0.316 0.263 0.300 0.264 0.302 0.254 0.297 0.247 0.311 0.246 0.314 0.245 0.306 96 0.141 0.194 0.141 0.194 0.141 0.196 - - - - - - 0.174 0.209 0.159 0.202 0.141 0.175 0.150 0.202 0.149 0.200 0.145 0.195 192 0.196 0.254 0.191 0.249 0.191 0.250 - - - - - - 0.231 0.258 0.210 0.249 0.186 0.220 0.192 0.243 0.194 0.245 0.188 0.240 336 0.260 0.309 0.251 0.302 0.249 0.301 - - - - - - 0.289 0.298 0.265 0.289 0.239 0.262 0.237 0.279 0.236 0.276 0.236 0.277 720 0.346 0.378 0.347 0.380 0.331 0.370 - - - - - - 0.376 0.347 0.362 0.344 0.319 0.315 0.314 0.335 0.304 0.326 0.313 0.335 Weather Avg. 0.236 0.284 0.233 0.281 0.228 0.279 - - - - - - 0.268 0.278 0.249 0.271 0.221 0.243 0.223 0.265 0.221 0.262 0.221 0.262 • * Weather included in the pre-training of TimesFM is not evaluated, denoted by a dash (–).
[figure]skip=0pt
Appendix D Limitations and Future Works
Our pruning metrics are calculated based on the training set, which may incur overfitting risks in extremely non-stationary time series data. To obtain a robust pruned model, possible solutions include data augmentation and adversarial attack techniques that keep the pruning metrics insensitive to noise. Furthermore, our pruning method works on TSFMs with strong zero-shot forecasting performance, since our basic assumption is that the prior knowledge of architecture specialization acquired by pre-training is effective for downstream forecasting. As for some tasks where pre-trained TSFMs show poor capabilities and cannot deliver an effective subnetwork, one possible solution is to apply pruning to a fine-tuned TSFM, rewind remaining parameters to the pre-trained weights, and fine-tune the pruned model. This follows prior works based on the classic Lottery Ticket Hypothesis [8], while we would like to leave it as a future work.

### v2

@@
 This “prune-then-finetune” paradigm often enables TSFMs to achieve state-of-the-art performance and surpass strong specialized baselines.
+Source code is made publicly available at https://github.com/SJTU-DMTai/Prune-then-Finetune.
 1 Introduction
@@
 Though only a few experts are activated for each token, all experts can be used and fine-tuned on the whole dataset.
-While some TSFMs claim superior zero-shot or full-shot performance over specialized models like PatchTST [23], their comparisons set the lookback length to 96 for specialized models while thousands for TSFMs.
+While some TSFMs claim state-of-the-art performance, their comparisons set the lookback length to 96 for specialized models (e.g., PatchTST [23]) while thousands for TSFMs.
 Li et al.
 demonstrate that specialized models with well-tuned hyperparameters outperform TSFMs, while they did not tune TSFM hyperparameters.
-In this work, after extensively tuning hyperparameters of all models based on the validation set, our benchmark results confirm that TSFMs, though equipped with specialization modules, still need fine-tuning for higher downstream performance.
-Also, we find that applying parameter-efficient fine-tuning [13] to TSFMs delivers similar performance improvement compared with full fine-tuning (see Appendix C.2).
+In this work, after extensively tuning hyperparameters of all models, our benchmark results confirm that TSFMs still cannot surpass specialized baselines.
+Also, we find that applying parameter-efficient fine-tuning [13] to TSFMs is not superior to full fine-tuning (see Appendix C.2) and cannot resolve the limitations in specialization.
+A concurrent work [24] also emphasizes that how to effectively fine-tuning TSFMs is an under-explored yet important problem.
 2.2 Model Pruning
@@
 The key statistics of the TSFMs are provided in Table 1.
-| Model | TTMA | Time-MoE | Time-MoE | Moirai | Moirai | Chronos | TimesFM | PatchTST |
-| Architecture | MLP | Decoder- Only | Decoder- Only | Encoder- Only | Encoder- Only | Encoder- Decoder | Decoder- Only | Encoder- Only |
-| Model Size | 5M | 113M | 453M | 91M | 311M | 205M | 500M | 500K |
-| #Layer | 6 | 12 | 12 | 12 | 24 | 24 | 50 | 3 |
-| #Head | \ | 12 | 12 | 12 | 16 | 12 | 16 | 12 |
-| 26384 | 384 | 768 | 768 | 1024 | 1024 | 1280 | 128 | |
+| Model | TTMA | Time-MoE | Time-MoE | Moirai | Moirai | Chronos | TimesFM | PatchTST | ||||||||||||||
+| Architecture | MLP |
+| Model Size | 5M | 113M | 453M | 91M | 311M | 205M | 500M | 500K | ||||||||||||||
+| #Layer | 6 | 12 | 12 | 12 | 24 | 24 | 50 | 3 | ||||||||||||||
+| #Head | \ | 12 | 12 | 12 | 16 | 12 | 16 | 12 | ||||||||||||||
+| 26384 | 384 | 768 | 768 | 1024 | 1024 | 1280 | 128 | |||||||||||||||
 | 52768 | 1536 | 3072 | 3072 | 4096 | 3072 | 1280 | 256 |

### v3

@@
 Time Series Foundation Models (TSFMs) are large forecasting models pre-trained on massive datasets.
-Recent researches on TSFMs [11, 26, 12, 2, 6, 29, 21] have demonstrated their impressive zero-shot forecasting capabilities across various domains without being specifically trained for those particular prediction tasks.
+Recent researches on TSFMs [garza2023timegpt1, Lag-Llama, goswami2024moment, ansari2024chronos, TimesFM, Moirai, Timer] have demonstrated their impressive zero-shot forecasting capabilities across various domains without being specifically trained for those particular prediction tasks.
 Time series forecasting has a distinct advantage compared to other machine learning tasks.
@@
 Our performance evaluations in Fig.
-1 demonstrate that even state-of-the-art pretrained TSFMs operating in zero-shot mode frequently underperform compared to specialized models like PatchTST [23].
+1 demonstrate that even state-of-the-art pretrained TSFMs operating in zero-shot mode frequently underperform compared to specialized models like PatchTST [Yuqietal-2023-PatchTST].
 This performance gap primarily stems from statistical differences between the diverse pre-training data and the specific patterns in downstream time series.
@@
 This approach risks overfitting and fails to capitalize on the valuable task-specific substructures identified during pre-training.
-In this paper, we propose decomposing TSFM deployment into two stages: architectural specialization through structured pruning [16, 4] and weight specialization through fine-tuning.
-While structure pruning by removing redundant components like channels or attention heads is commonly used to compress large models for lower latency and memory usage [22, 28], our primary objective is to facilitate fine-tuning in a narrowed parameter space for stable performance improvement.
-Specifically, for each downstream task, we analyze channel importance across all linear transformation layers of a TSFM using a loss-guided metric [22] to identify and remove unimportant model parameters.
+In this paper, we propose decomposing TSFM deployment into two stages: architectural specialization through structured pruning [li2017pruning, SurveyPruning] and weight specialization through fine-tuning.
+While structure pruning by removing redundant components like channels or attention heads is commonly used to compress large models for lower latency and memory usage [LLM-Pruner, wang2024modelcompressionefficientinference], our primary objective is to facilitate fine-tuning in a narrowed parameter space for stable performance improvement.
+Specifically, for each downstream task, we analyze channel importance across all linear transformation layers of a TSFM using a loss-guided metric [LLM-Pruner] to identify and remove unimportant model parameters.
 This results in a sparse model whose remaining parameters are fine-tuned on the downstream data.
@@
 Recently, increasing efforts have been devoted to pre-training TSFMs on large-scale time series data for a versatile starting point of various prediction tasks.
-Timer [21] and Timer-XL [20] [2] stack multiple Transformer layers with causal attention to capture temporal correlations.
+Timer [Timer] and Timer-XL [TimerXL] [ansari2024chronos] stack multiple Transformer layers with causal attention to capture temporal correlations.
 Furthermore, several specialization modules are designed to handle the heterogeneity of time series.
-For frequency-specific specialization, Moirai [29] differentiates input/output projection layers among frequencies, while TTM [7] and TimesFM [6] include categorical frequency embeddings as input.
-Time-MoE [27] and Moirai-MoE [19] adopt a sparse mixture-of-experts (MoE) design where tokens are processed by different expert feed-forward networks (FFNs).
+For frequency-specific specialization, Moirai [Moirai] differentiates input/output projection layers among frequencies, while TTM [ekambaram2024tinytimemixersttms] and TimesFM [TimesFM] include categorical frequency embeddings as input.
+Time-MoE [TimeMoE] and Moirai-MoE [Moirai-MoE] adopt a sparse mixture-of-experts (MoE) design where tokens are processed by different expert feed-forward networks (FFNs).
 Though only a few experts are activated for each token, all experts can be used and fine-tuned on the whole dataset.
-While some TSFMs claim state-of-the-art performance, their comparisons set the lookback length to 96 for specialized models (e.g., PatchTST [23]) while thousands for TSFMs.
-Li et al.
-demonstrate that specialized models with well-tuned hyperparameters outperform TSFMs, while they did not tune TSFM hyperparameters.
+While some TSFMs claim state-of-the-art performance, their comparisons set the lookback length to 96 for specialized models (e.g., PatchTST [Yuqietal-2023-PatchTST]) while thousands for TSFMs.
+li2024foundts demonstrate that specialized models with well-tuned hyperparameters outperform TSFMs, while they did not tune TSFM hyperparameters.
 In this work, after extensively tuning hyperparameters of all models, our benchmark results confirm that TSFMs still cannot surpass specialized baselines.
-Also, we find that applying parameter-efficient fine-tuning [13] to TSFMs is not superior to full fine-tuning (see Appendix C.2) and cannot resolve the limitations in specialization.
-A concurrent work [24] also emphasizes that how to effectively fine-tuning TSFMs is an under-explored yet important problem.
+Also, we find that applying parameter-efficient fine-tuning [LoRA] to TSFMs is not superior to full fine-tuning (see Appendix C.3) and cannot resolve the limitations in specialization.
+A concurrent work [qiao2025multiscalefinetuningencoderbasedtime] also emphasizes that how to effectively fine-tuning TSFMs is an under-explored yet important problem.
 2.2 Model Pruning
-Pruning [5], a conventional technique for network compression and acceleration, is increasingly applied to Large Language Models (LLMs) for efficient inference.
+Pruning [Optimal-brain-damage], a conventional technique for network compression and acceleration, is increasingly applied to Large Language Models (LLMs) for efficient inference.
 Based on pruning units, pruning has two categories.
-Unstructured pruning removes unimportant elements of weight matrices, while the sparse matrices require specialized hardware or software optimizations for acceleration [35].
+Unstructured pruning removes unimportant elements of weight matrices, while the sparse matrices require specialized hardware or software optimizations for acceleration [zhu-etal-2024-survey-model-compression].
 Structured pruning removes entire components (e.g., channels and attention heads with the lowest importance) and is hardware-friendly.
 Numerous pruning metrics have been proposed to determine whether a pruning unit is essential and should be pruned or not.
-Various metrics guide pruning; for example, LLM-Pruner [22] assesses a component’s importance by its impact on loss.
-Typically, LLMs after structure pruning struggle to regain original performance even after retraining [22, 3, 35].
+Various metrics guide pruning; for example, LLM-Pruner [LLM-Pruner] assesses a component’s importance by its impact on loss.
+Typically, LLMs after structure pruning struggle to regain original performance even after retraining [LLM-Pruner, ashkboos2024slicegpt, zhu-etal-2024-survey-model-compression].
 The potential reason behind lossy compression of LLMs is that various neurons can preserve language skills and vast factual knowledge, which is critical to commonsense questions.
@@
 Instead, we are devoted to improving the fine-tuning performance of TSFMs by pruning neurons that are less relevant to the downstream task.
-The rationale is that time series are characterized by strong heterogeneity across domains and frequencies [19], and a specific task may only need to activate a certain fraction of model parameters [27].
+The rationale is that time series are characterized by strong heterogeneity across domains and frequencies [Moirai-MoE], and a specific task may only need to activate a certain fraction of model parameters [TimeMoE].
 Moreover, due to the nature of non-stationarity, time series usually exhibit various data patterns over time, and there are scarce latent features that maintain long-term effectiveness for time series forecasting.
@@
 where is the activation function (e.g., ReLU, GELU, SwiGLU) that can deactivate some intermediate hidden states as zeros or tiny negative values.
-TTM [7], though not based on Transformers, also incorporates MLPs with GELU activation.
+TTM [ekambaram2024tinytimemixersttms], though not based on Transformers, also incorporates MLPs with GELU activation.
 For the -th intermediate FFN channel, the probability of producing positive activations over a dataset can reflect its relevance to the downstream task.
@@
 3.2 Empirical Analysis of TSFMs
-According to the theory of bias-variance trade-offs [9], a large model can induce high variance risks in predictions, which seems contradictory to the skilled zero-shot performance of TSFMs on unseen tasks.
+According to the theory of bias-variance trade-offs [friedman2001elements], a large model can induce high variance risks in predictions, which seems contradictory to the skilled zero-shot performance of TSFMs on unseen tasks.
 To ascertain the mechanism of TSFMs, we conduct an investigation into TSFMs through the lens of attention heads and activations in FFNs.
-We perform 96-step forecasting on the Weather [30] and ETTm1 [34] datasets by PatchTST and seven pre-trained TSFMs, including TTM [7], Time-MoE, Time-MoE [27], Moirai, Moirai [29], Chronos-bolt [2], and TimesFM-2.0-500M [6].
+We perform 96-step forecasting on the Weather [Autoformer] and ETTm1 [zhou2021informer] datasets by PatchTST and seven pre-trained TSFMs, including TTM [ekambaram2024tinytimemixersttms], Time-MoE, Time-MoE [TimeMoE], Moirai, Moirai [Moirai], Chronos-bolt [ansari2024chronos], and TimesFM-2.0-500M [TimesFM].
 The key statistics of the TSFMs are provided in Table 1.
-| Model | TTMA | Time-MoE | Time-MoE | Moirai | Moirai | Chronos | TimesFM | PatchTST | ||||||||||||||
-| Architecture | MLP |
-| Model Size | 5M | 113M | 453M | 91M | 311M | 205M | 500M | 500K | ||||||||||||||
-| #Layer | 6 | 12 | 12 | 12 | 24 | 24 | 50 | 3 | ||||||||||||||
-| #Head | \ | 12 | 12 | 12 | 16 | 12 | 16 | 12 | ||||||||||||||
-| 26384 | 384 | 768 | 768 | 1024 | 1024 | 1280 | 128 | |||||||||||||||
+| Model | TTMA | Time-MoE | Time-MoE | Moirai | Moirai | Chronos | TimesFM | PatchTST |
+| Architecture | MLP | Decoder- Only | Decoder- Only | Encoder- Only | Encoder- Only | Encoder- Decoder | Decoder- Only | Encoder- Only |
+| Model Size | 5M | 113M | 453M | 91M | 311M | 205M | 500M | 500K |
+| #Layer | 6 | 12 | 12 | 12 | 24 | 24 | 50 | 3 |
+| #Head | \ | 12 | 12 | 12 | 16 | 12 | 16 | 12 |
+| 26384 | 384 | 768 | 768 | 1024 | 1024 | 1280 | 128 | |
 | 52768 | 1536 | 3072 | 3072 | 4096 | 3072 | 1280 | 256 |
@@
 3.2.2 Sparse Activations in FFNs
-Shi et al.
-and Liu et al.
-found that MoE-based TSFMs tend to allocate different expert FFNs for samples from diverse datasets, and sparsely activated experts can address the heterogeneity of time series.
+TimeMoE and Moirai-MoE found that MoE-based TSFMs tend to allocate different expert FFNs for samples from diverse datasets, and sparsely activated experts can address the heterogeneity of time series.
 We assume that there is also a sparse activation pattern at the granularity of FFN channels.
@@
 4.2 Importance Score of Pruning Unit
-From the perspective of forecasting loss, a redundant channel, if pruned, would result in negligible loss changes. Following prior works [22], we define the channel-wise importance score as follows:
+From the perspective of forecasting loss, a redundant channel, if pruned, would result in negligible loss changes.
+Following prior works [LLM-Pruner], we define the channel-wise importance score as follows:
 Definition 3
@@
 In Eq.
-(6), we employ Fisher information to approximate the second-order derivative by the square of the first-order derivative [18].
+(6), we employ Fisher information to approximate the second-order derivative by the square of the first-order derivative [GroupFisher].
 The derivative w.r.t.
@@
 This channel redundancy is notably more pronounced in larger TSFM variants.
-This TSFM behavior caters to the nature of time series forecasting, where the presence of concept drift [10, 14, 31, 33, 32] implies that enduringly effective data patterns are scarce.
+This TSFM behavior caters to the nature of temporal data, where the presence of concept drift [concept_drift, kim2021reversible, DoubleAdapt, Proceed, LIFT, duan2025brain] implies that enduringly effective data patterns are scarce.
 Consequently, structured pruning emerges as a compelling strategy to selectively restrain the number of exploited latent features.
@@
 We compare full-shot forecasting performance with and without pruning, analyze the resulting sparsity patterns and inference speedups, and assess the transferability of the pruned TSFMs to other datasets.
-Additional experimental results are provided in Appendix C, including ablation studies on the tuning methods (Table 7) and the pruning methods (Table 8).
+Additional experimental results are provided in Appendix C, including short-term forecasting results on the M4 benchmark [M4] and ablation studies on the tuning methods (Table 8) and the pruning methods (Table 9).
 5.1 Experimental Settings
@@
 Datasets.
-We evaluate long-term forecasting performance on three widely used benchmarks: ETT [34], Weather [30], and Electricity [15].
+We evaluate long-term forecasting performance on three widely used benchmarks: ETT [zhou2021informer], Weather [Autoformer], and Electricity [LSTNet].
 Dataset statistics and train/validation/test splits are summarized in Table 5.
@@
 Different from random dropout, pruning focuses on a task-relevant subspace of parameters consistently across fine-tuning and inference.
-Model TimeMoE TimeMoE Timer-XL Moirai Moirai Chronos TTMA PatchTST Method FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT Tasks MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE 96 0.340 0.377 0.340 0.376 0.331 0.372 0.329 0.372 0.365 0.393 0.353 0.391 0.367 0.396 0.366 0.396 0.371 0.391 0.370 0.391 0.366 0.374 0.366 0.375 0.359 0.395 0.358 0.393 0.376 0.396 192 0.380 0.408 0.376 0.404 0.371 0.398 0.369 0.399 0.400 0.416 0.389 0.413 0.405 0.418 0.404 0.418 0.416 0.416 0.413 0.416 0.418 0.405 0.417 0.405 0.392 0.417 0.391 0.410 0.399 0.416 336 0.406 0.432 0.399 0.425 0.392 0.413 0.389 0.411 0.419 0.431 0.409 0.426 0.435 0.434 0.429 0.430 0.455 0.431 0.455 0.432 0.453 0.424 0.449 0.422 0.407 0.430 0.402 0.424 0.418 0.432 720 0.437 0.469 0.424 0.457 0.403 0.438 0.387 0.419 0.454 0.465 0.433 0.455 0.465 0.459 0.427 0.441 0.475 0.457 0.454 0.444 0.452 0.441 0.451 0.437 0.448 0.467 0.449 0.465 0.450 0.469 Avg 0.391 0.422 0.385 0.416 0.374 0.405 0.369 0.400 0.410 0.426 0.396 0.421 0.418 0.427 0.407 0.421 0.429 0.424 0.423 0.421 0.422 0.411 0.421 0.410 0.402 0.427 0.400 0.423 0.411 0.428 ETTh1 #p 63% 63% 67% 58%69% 36%39% 55% 91%93% 96 0.267 0.330 0.262 0.327 0.272 0.332 0.272 0.332 0.287 0.345 0.275 0.337 0.273 0.338 0.272 0.338 0.288 0.341 0.287 0.341 0.266 0.311 0.266 0.311 0.264 0.333 0.262 0.328 0.277 0.339 192 0.341 0.380 0.331 0.374 0.356 0.383 0.356 0.382 0.344 0.385 0.334 0.378 0.331 0.385 0.331 0.385 0.349 0.386 0.350 0.387 0.334 0.357 0.331 0.356 0.321 0.372 0.324 0.370 0.345 0.381 336 0.386 0.415 0.372 0.403 0.420 0.424 0.411 0.418 0.375 0.410 0.368 0.404 0.358 0.410 0.358 0.409 0.373 0.409 0.371 0.408 0.378 0.391 0.370 0.388 0.351 0.397 0.351 0.396 0.368 0.404 720 0.419 0.454 0.377 0.423 0.534 0.494 0.459 0.454 0.411 0.445 0.398 0.439 0.392 0.444 0.388 0.442 0.402 0.439 0.394 0.435 0.395 0.418 0.389 0.415 0.395 0.439 0.393 0.437 0.397 0.432 Avg 0.353 0.395 0.336 0.382 0.396 0.408 0.375 0.397 0.354 0.396 0.344 0.390 0.339 0.394 0.337 0.394 0.353 0.394 0.351 0.393 0.343 0.369 0.339 0.368 0.333 0.385 0.333 0.383 0.347 0.389 ETTh2 #p 82% 77% 68% 55%78% 27%36% 56% 93%98% 96 0.220 0.302 0.220 0.302 0.214 0.305 0.211 0.299 0.305 0.355 0.297 0.354 0.309 0.343 0.304 0.342 0.311 0.345 0.312 0.346 0.285 0.315 0.283 0.315 0.295 0.346 0.291 0.340 0.290 0.343 192 0.284 0.347 0.282 0.344 0.268 0.345 0.257 0.333 0.348 0.383 0.337 0.381 0.349 0.367 0.341 0.366 0.343 0.370 0.342 0.370 0.329 0.345 0.326 0.344 0.332 0.366 0.327 0.363 0.329 0.368 336 0.349 0.392 0.335 0.381 0.326 0.386 0.300 0.364 0.382 0.406 0.367 0.403 0.379 0.385 0.375 0.384 0.364 0.385 0.359 0.385 0.361 0.369 0.356 0.367 0.356 0.383 0.356 0.381 0.360 0.390 720 0.462 0.466 0.415 0.437 0.428 0.449 0.380 0.415 0.428 0.439 0.409 0.434 0.410 0.409 0.397 0.409 0.401 0.409 0.400 0.409 0.434 0.413 0.423 0.407 0.409 0.420 0.395 0.406 0.416 0.422 Avg 0.329 0.377 0.313 0.366 0.309 0.371 0.287 0.353 0.366 0.396 0.353 0.393 0.362 0.376 0.354 0.375 0.355 0.377 0.353 0.378 0.352 0.361 0.347 0.358 0.348 0.379 0.342 0.373 0.349 0.381 ETTm1 #p 92% 64% 81% 51%60% 51%52% 55% 86%99% 96 0.161 0.253 0.158 0.251 0.166 0.258 0.164 0.258 0.191 0.280 0.181 0.273 0.174 0.261 0.172 0.259 0.165 0.251 0.165 0.251 0.160 0.231 0.157 0.231 0.160 0.251 0.160 0.248 0.165 0.254 192 0.229 0.308 0.217 0.299 0.248 0.318 0.236 0.313 0.245 0.320 0.233 0.310 0.234 0.304 0.231 0.300 0.234 0.304 0.232 0.301 0.225 0.276 0.219 0.274 0.213 0.289 0.213 0.286 0.221 0.292 336 0.313 0.367 0.273 0.342 0.347 0.381 0.314 0.367 0.297 0.356 0.283 0.344 0.284 0.340 0.282 0.337 0.288 0.343 0.287 0.340 0.286 0.317 0.276 0.312 0.269 0.326 0.268 0.322 0.275 0.325 720 0.486 0.464 0.375 0.407 0.552 0.488 0.471 0.453 0.389 0.416 0.367 0.396 0.353 0.386 0.350 0.384 0.366 0.397 0.365 0.396 0.381 0.377 0.363 0.369 0.346 0.379 0.338 0.369 0.360 0.380 Avg 0.297 0.348 0.256 0.325 0.328 0.361 0.296 0.348 0.281 0.343 0.266 0.331 0.261 0.323 0.259 0.320 0.263 0.324 0.262 0.322 0.263 0.300 0.254 0.297 0.247 0.311 0.245 0.306 0.255 0.313 ETTm2 #p 75% 57% 64% 51%56% 35%38% 55% 86%99% 96 0.141 0.194 0.141 0.196 0.140 0.193 0.139 0.190 0.169 0.226 0.165 0.219 0.178 0.228 0.159 0.222 0.147 0.199 0.147 0.198 0.174 0.209 0.141 0.175 0.150 0.202 0.145 0.195 0.149 0.196 192 0.196 0.254 0.191 0.250 0.191 0.245 0.189 0.241 0.217 0.271 0.209 0.261 0.220 0.269 0.216 0.265 0.211 0.263 0.198 0.255 0.231 0.258 0.186 0.220 0.192 0.243 0.188 0.240 0.193 0.240 336 0.260 0.309 0.249 0.301 0.255 0.297 0.253 0.292 0.267 0.308 0.255 0.296 0.308 0.330 0.238 0.287 0.277 0.315 0.279 0.317 0.289 0.298 0.239 0.262 0.237 0.279 0.236 0.277 0.244 0.281 720 0.346 0.378 0.331 0.370 0.371 0.383 0.361 0.370 0.336 0.358 0.322 0.345 0.390 0.388 0.333 0.361 0.374 0.383 0.345 0.365 0.376 0.347 0.319 0.315 0.314 0.335 0.313 0.335 0.314 0.332 Avg 0.236 0.284 0.228 0.279 0.239 0.280 0.236 0.273 0.247 0.291 0.238 0.280 0.274 0.304 0.237 0.284 0.252 0.290 0.242 0.284 0.268 0.278 0.221 0.243 0.223 0.265 0.221 0.262 0.225 0.262 Weather #p 68% 88% 75% 20%62% 20%44% 55% 91%94% 96 - - - - - - - - 0.139 0.237 0.138 0.237 0.152 0.247 0.135 0.236 0.148 0.243 0.138 0.236 0.121 0.208 0.121 0.208 0.129 0.226 0.129 0.226 0.133 0.233 192 - - - - - - - - 0.158 0.256 0.157 0.255 0.170 0.263 0.162 0.263 0.172 0.265 0.164 0.265 0.139 0.226 0.139 0.225 0.147 0.243 0.148 0.244 0.150 0.248 336 - - - - - - - - 0.177 0.275 0.176 0.275 0.192 0.285 0.176 0.275 0.187 0.280 0.190 0.282 0.157 0.244 0.156 0.243 0.167 0.266 0.166 0.265 0.168 0.267 720 - - - - - - - - 0.219 0.313 0.217 0.312 0.220 0.307 0.210 0.305 0.219 0.309 0.215 0.310 0.199 0.283 0.195 0.278 0.202 0.297 0.199 0.295 0.202 0.295 Avg - - - - - - - - 0.173 0.270 0.172 0.270 0.184 0.276 0.171 0.270 0.182 0.274 0.177 0.273 0.154 0.240 0.153 0.239 0.161 0.258 0.161 0.258 0.163 0.261 Electricity #p 56% 40%42% 3%60% 53% 94%98% Win Rate 46.0% 68.0% 46.0% 52.0% 11.7% 26.7% 25.0% 33.3% 18.3% 20.0% 80.0% 95.0% 85.0% 93.3% • * “#p” is the proportion of the remaining parameters after pruning relative to the original ones.
+Model TimeMoE TimeMoE Timer-XL Moirai Moirai Chronos TTMA PatchTST Method FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT FT Prune+FT Tasks MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE 96 0.340 0.377 0.340 0.376 0.331 0.372 0.329 0.372 0.365 0.393 0.353 0.391 0.367 0.396 0.366 0.396 0.371 0.391 0.370 0.391 0.366 0.374 0.366 0.375 0.359 0.395 0.358 0.393 0.376 0.396 192 0.380 0.408 0.376 0.404 0.371 0.398 0.369 0.399 0.400 0.416 0.389 0.413 0.405 0.418 0.404 0.418 0.416 0.416 0.413 0.416 0.418 0.405 0.417 0.405 0.392 0.417 0.391 0.410 0.399 0.416 336 0.406 0.432 0.399 0.425 0.392 0.413 0.389 0.411 0.419 0.431 0.409 0.426 0.435 0.434 0.429 0.430 0.455 0.431 0.455 0.432 0.453 0.424 0.449 0.422 0.407 0.430 0.402 0.424 0.418 0.432 720 0.437 0.469 0.424 0.457 0.403 0.438 0.387 0.419 0.454 0.465 0.433 0.455 0.465 0.459 0.427 0.441 0.475 0.457 0.454 0.444 0.452 0.441 0.451 0.437 0.448 0.467 0.449 0.465 0.450 0.469 Avg 0.391 0.422 0.385 0.416 0.374 0.405 0.369 0.400 0.410 0.426 0.396 0.421 0.418 0.427 0.407 0.421 0.429 0.424 0.423 0.421 0.422 0.411 0.421 0.410 0.402 0.427 0.400 0.423 0.411 0.428 ETTh1 #p 63% 63% 67% 58%69% 36%39% 55% 91%93% 96 0.267 0.330 0.262 0.327 0.272 0.332 0.272 0.332 0.287 0.345 0.275 0.337 0.273 0.338 0.272 0.338 0.288 0.341 0.287 0.341 0.266 0.311 0.266 0.311 0.264 0.333 0.262 0.328 0.277 0.339 192 0.341 0.380 0.331 0.374 0.356 0.383 0.356 0.382 0.344 0.385 0.334 0.378 0.331 0.385 0.331 0.385 0.349 0.386 0.350 0.387 0.334 0.357 0.331 0.356 0.321 0.372 0.324 0.370 0.345 0.381 336 0.386 0.415 0.372 0.403 0.420 0.424 0.411 0.418 0.375 0.410 0.368 0.404 0.358 0.410 0.358 0.409 0.373 0.409 0.371 0.408 0.378 0.391 0.370 0.388 0.351 0.397 0.351 0.396 0.368 0.404 720 0.419 0.454 0.377 0.423 0.534 0.494 0.459 0.454 0.411 0.445 0.398 0.439 0.392 0.444 0.388 0.442 0.402 0.439 0.394 0.435 0.395 0.418 0.389 0.415 0.395 0.439 0.393 0.437 0.397 0.432 Avg 0.353 0.395 0.336 0.382 0.396 0.408 0.375 0.397 0.354 0.396 0.344 0.390 0.339 0.394 0.337 0.394 0.353 0.394 0.351 0.393 0.343 0.369 0.339 0.368 0.333 0.385 0.333 0.383 0.347 0.389 ETTh2 #p 82% 77% 68% 55%78% 27%36% 56% 93%98% 96 0.220 0.302 0.220 0.302 0.214 0.305 0.211 0.299 0.305 0.355 0.297 0.354 0.309 0.343 0.304 0.342 0.311 0.345 0.312 0.346 0.285 0.315 0.282 0.315 0.295 0.346 0.291 0.340 0.290 0.343 192 0.284 0.347 0.282 0.344 0.268 0.345 0.257 0.333 0.348 0.383 0.337 0.381 0.349 0.367 0.341 0.366 0.343 0.370 0.342 0.370 0.329 0.345 0.323 0.343 0.332 0.366 0.327 0.363 0.329 0.368 336 0.349 0.392 0.335 0.381 0.326 0.386 0.300 0.364 0.382 0.406 0.367 0.403 0.379 0.385 0.375 0.384 0.364 0.385 0.359 0.385 0.361 0.369 0.350 0.365 0.356 0.383 0.356 0.381 0.360 0.390 720 0.462 0.466 0.415 0.437 0.428 0.449 0.380 0.415 0.428 0.439 0.409 0.434 0.410 0.409 0.397 0.409 0.401 0.409 0.400 0.409 0.434 0.413 0.412 0.404 0.409 0.420 0.395 0.406 0.416 0.422 Avg.
+0.329 0.377 0.313 0.366 0.309 0.371 0.287 0.353 0.366 0.396 0.353 0.393 0.362 0.376 0.354 0.375 0.355 0.377 0.353 0.378 0.352 0.361 0.342 0.357 0.348 0.379 0.342 0.373 0.349 0.381 ETTm1 #p 92% 64% 81% 51%60% 51%52% 18% 86%99% 96 0.161 0.253 0.158 0.251 0.166 0.258 0.164 0.258 0.191 0.280 0.181 0.273 0.174 0.261 0.172 0.259 0.165 0.251 0.165 0.251 0.160 0.231 0.156 0.234 0.160 0.251 0.160 0.248 0.165 0.254 192 0.229 0.308 0.217 0.299 0.248 0.318 0.236 0.313 0.245 0.320 0.233 0.310 0.234 0.304 0.231 0.300 0.234 0.304 0.232 0.301 0.225 0.276 0.215 0.275 0.213 0.289 0.213 0.286 0.221 0.292 336 0.313 0.367 0.273 0.342 0.347 0.381 0.314 0.367 0.297 0.356 0.283 0.344 0.284 0.340 0.282 0.337 0.288 0.343 0.287 0.340 0.286 0.317 0.268 0.313 0.269 0.326 0.268 0.322 0.275 0.325 720 0.486 0.464 0.375 0.407 0.552 0.488 0.471 0.453 0.389 0.416 0.367 0.396 0.353 0.386 0.350 0.384 0.366 0.397 0.365 0.396 0.381 0.377 0.348 0.367 0.346 0.379 0.338 0.369 0.360 0.380 Avg.
+0.297 0.348 0.256 0.325 0.328 0.361 0.296 0.348 0.281 0.343 0.266 0.331 0.261 0.323 0.259 0.320 0.263 0.324 0.262 0.322 0.263 0.300 0.247 0.297 0.247 0.311 0.245 0.306 0.255 0.313 ETTm2 #p 75% 57% 64% 51%56% 35%38% 18% 86%99% 96 0.141 0.194 0.141 0.196 0.140 0.193 0.139 0.190 0.169 0.226 0.165 0.219 0.178 0.228 0.159 0.222 0.147 0.199 0.147 0.198 0.174 0.209 0.143 0.180 0.150 0.202 0.145 0.195 0.149 0.196 192 0.196 0.254 0.191 0.250 0.191 0.245 0.189 0.241 0.217 0.271 0.209 0.261 0.220 0.269 0.216 0.265 0.211 0.263 0.198 0.255 0.231 0.258 0.186 0.223 0.192 0.243 0.188 0.240 0.193 0.240 336 0.260 0.309 0.249 0.301 0.255 0.297 0.253 0.292 0.267 0.308 0.255 0.296 0.308 0.330 0.238 0.287 0.277 0.315 0.279 0.317 0.289 0.298 0.235 0.262 0.237 0.279 0.236 0.277 0.244 0.281 720 0.346 0.378 0.331 0.370 0.371 0.383 0.361 0.370 0.336 0.358 0.322 0.345 0.390 0.388 0.333 0.361 0.374 0.383 0.345 0.365 0.376 0.347 0.296 0.306 0.314 0.335 0.313 0.335 0.314 0.332 Avg.
+0.236 0.284 0.228 0.279 0.239 0.280 0.236 0.273 0.247 0.291 0.238 0.280 0.274 0.304 0.237 0.284 0.252 0.290 0.242 0.284 0.268 0.278 0.215 0.243 0.223 0.265 0.221 0.262 0.225 0.262 Weather #p 68% 88% 75% 20%62% 20%44% 17% 91%94% 96 - - - - - - - - 0.139 0.237 0.138 0.237 0.152 0.247 0.135 0.236 0.148 0.243 0.138 0.236 0.121 0.208 0.120 0.206 0.129 0.226 0.129 0.226 0.133 0.233 192 - - - - - - - - 0.158 0.256 0.157 0.255 0.170 0.263 0.162 0.263 0.172 0.265 0.164 0.265 0.139 0.226 0.138 0.224 0.147 0.243 0.148 0.243 0.150 0.248 336 - - - - - - - - 0.177 0.275 0.176 0.275 0.192 0.285 0.176 0.275 0.187 0.280 0.175 0.283 0.157 0.244 0.155 0.242 0.167 0.266 0.165 0.263 0.168 0.267 720 - - - - - - - - 0.219 0.313 0.217 0.312 0.220 0.307 0.210 0.305 0.219 0.309 0.215 0.310 0.199 0.283 0.194 0.277 0.202 0.297 0.199 0.293 0.202 0.295 Avg.
+- - - - - - - - 0.173 0.270 0.172 0.270 0.184 0.276 0.171 0.270 0.182 0.274 0.173 0.274 0.154 0.240 0.152 0.237 0.161 0.258 0.160 0.256 0.163 0.261 Electricity #p 56% 40%42% 3%60% 53% 94%98% Win Rate 46.0% 68.0% 46.0% 52.0% 11.7% 26.7% 25.0% 33.3% 18.3% 20.0% 80.0% 98.3% 85.0% 93.3% • * “#p” is the proportion of the remaining parameters after pruning relative to the original ones.
 • * “Win Rate” is based on comparisons between each TSFM and PatchTST over all datasets.
 • * Electricity included in the pre-training of Time-MoE is not evaluated, denoted by a dash (–).
-• * We present performance comparison about TimesFM in Table 8.
+• * We present performance comparison about TimesFM and the Traffic [Autoformer] dataset in Table 9.
 5.3 Sparsification Analysis
@@
 Although both models consist of the same number of Transformer layers, Moirai adopts an encoder-only architecture, while Chronos follows an encoder-decoder design.
-Moirai generally exhibits more aggressive pruning in its deeper layers, with non-negligible pruning also observed in its shallow layers.
-In contrast, pruning in Chronos is predominantly concentrated in the decoder (i.e., the last 12 layers), while the encoder remains largely intact.
+Both models generally exhibits more aggressive pruning in their deeper layers, with non-negligible pruning also observed in the shallow layers.
+In contrast, pruning in Chronos is more concentrated in the decoder (i.e., the last 12 layers) than the encoder.
 We attribute this to its use of a much smaller patch size for time series tokenization, which requires the encoder to capture and aggregate more contextual information across tokens.
@@
 | Moirai | 1.86 | 0.25 | 7.4 |
-| Chronos | 8.51 | 7.19 | 1.2 |
+| Chronos | 8.51 | 1.26 | 6.5 |
 | TTMA | 0.01 | 0.01 | 1.0 |
@@
 This work pioneers structured pruning primarily for performance enhancement via architectural specialization, offering a promising pathway to unlock the full potential of TSFMs in practical applications.
+Acknowledgments and Disclosure of Funding
+This work is supported by the National Key Research and Development Program of China (No.
+2023YFB4502902), National Natural Science Foundation of China (No.
+62522211), and Key Research and Development Program of Xinjiang Uygur Autonomous Region (Grant No.
+2023B01027, 2023B01027-1).
+NeurIPS Paper Checklist
+1.
+Claims
+Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+Answer: [Yes]
+Guidelines:
+The answer NA means that the abstract and introduction do not include the claims made in the paper.
+The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+2.
+Limitations
+Question: Does the paper discuss the limitations of the work performed by the authors?
+Answer: [Yes]
+Guidelines:
+The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+The authors are encouraged to create a separate "Limitations" section in their paper.
+The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+3.
+Theory assumptions and proofs
+Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+Answer: [N/A]
+Justification: the paper does not include theoretical results.
+Guidelines:
+The answer NA means that the paper does not include theoretical results.
+All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+All assumptions should be clearly stated or referenced in the statement of any theorems.
+The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+Theorems and Lemmas that the proof relies upon should be properly referenced.
+4.
+Experimental result reproducibility
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+Answer: [Yes]
+Guidelines:
+The answer NA means that the paper does not include experiments.
+If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+5.
+Open access to data and code
+Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+Answer: [Yes]
+Guidelines:
+The answer NA means that paper does not include experiments requiring code.
+Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+6.
+Experimental setting/details
+Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+Answer: [Yes]
+Guidelines:
+The answer NA means that the paper does not include experiments.
+The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+The full details can be provided either with the code, in appendix, or as supplemental material.
+7.
+Experiment statistical significance
+Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+Answer: [No]
+Justification: The computational cost is unaffordable to repeat extensive experiments on 764 across seven large models, six datasets, and four prediction horizons.
+Guidelines:
+The answer NA means that the paper does not include experiments.
+The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+The assumptions made should be given (e.g., Normally distributed errors).
+It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+8.
+Experiments compute resources
+Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+Answer: [Yes]
+Guidelines:
+The answer NA means that the paper does not include experiments.
+The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+9.
+Code of ethics
+Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+Answer: [Yes]
+Guidelines:
+The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+10.
+Broader impacts
+Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+Answer: [N/A]
+Justification: there is no societal impact of the work performed.
+Guidelines:
+The answer NA means that there is no societal impact of the work performed.
+If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+11.
+Safeguards
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+Answer: [N/A]
+Justification: the paper poses no such risks.
+Guidelines:
+The answer NA means that the paper poses no such risks.
+Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+12.
+Licenses for existing assets
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+Answer: [Yes]
+Justification: Yes
+Guidelines:
+The answer NA means that the paper does not use existing assets.
+The authors should cite the original paper that produced the code package or dataset.
+The authors should state which version of the asset is used and, if possible, include a URL.
+The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+13.
+New assets
+Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+Answer: [N/A]
+Justification: the paper does not release new assets.
+Guidelines:
+The answer NA means that the paper does not release new assets.
+Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+The paper should discuss whether and how consent was obtained from people whose asset is used.
+At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+14.
+Crowdsourcing and research with human subjects
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+Answer: [N/A]
+Justification: does not involve crowdsourcing nor research with human subjects.
+Guidelines:
+The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+15.
+Institutional review board (IRB) approvals or equivalent for research with human subjects
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+Answer: [N/A]
+Justification: does not involve crowdsourcing nor research with human subjects
+Guidelines:
+The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+16.
+Declaration of LLM usage
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+Answer: [N/A]
+Justification: Only for writing
+Guidelines:
+The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix A Implementation Details
@@
 | Electricity | 321 | (12280, 1755, 3509) | Hourly | electricity load |
-We evaluate the long-term forecasting performance across six well-established datasets, including the Weather [30], Electricity [15], and ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) [34].
+We evaluate the long-term forecasting performance across six well-established datasets, including the Weather [Autoformer], Electricity [LSTNet], and ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) [zhou2021informer].
 A detailed description of each dataset is provided in Table 5.
-We use the full test set without the “drop-last” strategy used in PatchTST [23].
-We borrow benchmark results of PatchTST from TFB [25] that carefully tuned the hyperparameters.
+We use the full test set without the “drop-last” strategy used in PatchTST [Yuqietal-2023-PatchTST].
+We borrow benchmark results of PatchTST from TFB [qiu2024tfb] that carefully tuned the hyperparameters.
 A.2 Experimental Details
@@
 Appendix C Additional Experimental Results
-C.1 Sparsification Visualization
+C.1 Short-term Forecasting Performance
+Table 7 compares short-term forecasting performance of Chronos with different tuning methods over the M4 benchmark [M4].
+Better performance can be indicated by smaller evaluation metrics including MSE (mean-square error), MAE (mean absolute error), MASE (mean absolute scaled error), CRPS (continuous ranked probability score), and NRMSE (normalized root-mean-square error), where MSE and MAE are calculated based on raw values rather than normalized ones.
+In most cases, our proposed method can lead to better short-term forecasting performance than the naive full fine-tuning method.
+| Chronos | MSE | MAE | MASE | CRPS | NRMSE | |
+| Yearly (=6) | Zero-shot | 3926423 | 943 | 3.507 | 0.121 | 0.318 |
+| Fine-Tune | 3226941 | 833 | 3.089 | 0.107 | 0.309 | |
+| Prune+FT | 3125398 | 816 | 3.007 | 0.104 | 0.283 | |
+| Quarterly (=8) | Zero-shot | 1878597 | 578 | 1.224 | 0.077 | 0.229 |
+| Fine-Tune | 1775734 | 571 | 1.221 | 0.075 | 0.223 | |
+| Prune+FT | 1749123 | 551 | 1.225 | 0.074 | 0.221 | |
+| Monthly (=18) | Zero-shot | 1934416 | 565 | 0.949 | 0.094 | 0.289 |
+| Fine-Tune | 1894232 | 571 | 0.946 | 0.095 | 0.286 | |
+| Prune+FT | 1820581 | 556 | 0.920 | 0.091 | 0.280 | |
+| Weekly (=13) | Zero-shot | 240658 | 258 | 2.08 | 0.038 | 0.089 |
+| Fine-Tune | 239732 | 257 | 2.07 | 0.038 | 0.089 | |
+| Prune+FT | 238142 | 253 | 1.99 | 0.037 | 0.089 | |
+| Daily (=14) | Zero-shot | 359007 | 172 | 3.20 | 0.093 | 0.021 |
+| Fine-Tune | 511617 | 175 | 3.21 | 0.110 | 0.021 | |
+| Prune+FT | 536501 | 168 | 3.05 | 0.113 | 0.021 | |
+| Hourly (=48) | Zero-shot | 1154587 | 244 | 0.837 | 0.025 | 0.147 |
+| Fine-Tune | 1032867 | 239 | 0.943 | 0.026 | 0.139 | |
+| Prune+FT | 937494 | 227 | 0.805 | 0.024 | 0.131 |
+[figure]skip=0pt
+C.2 Sparsification Visualization
 We visualize more sparsification results in Fig.
-10.
+9.
 Moirai and Timer-XL exhibit sparsification in input channels of output projections and intermediate channels of FFNs, which confirms that pruning based solely on importance scores can handle the inherent sparsity of TSFMs.
@@
 In our supplementary materials, we will provide results of pruning the attention heads based on a threshold of the relative output norm.
-C.2 Performance of Other Tuning Methods
+C.3 Performance of Other Tuning Methods
 | Models | Chronos | TTMA | |||||||||||||||
@@
 Parameter-Efficient Fine-Tuning (PEFT) has become a popular approach to efficient LLM fine-tuning by reducing the number of trainable parameters.
-LoRA [13] is one of the most prevalent PEFT methods that learns two additional low-rank weight matrices as an adjustment to the original full-rank weight matrix at each layer.
+LoRA [LoRA] is one of the most prevalent PEFT methods that learns two additional low-rank weight matrices as an adjustment to the original full-rank weight matrix at each layer.
 In this section, we apply LoRA to Chronos and TTMA with the rank hyperparameter chosen from {32, 64, 128}.
@@
 Thus, we would like to adopt full fine-tuning for better performance.
-C.3 Performance of Variant Pruning Methods
+C.4 Performance of Variant Pruning Methods
 In light of the inherent sparsity of TSFMs, we introduce a variant of TSFM pruning, denoted as Prune.
@@
 0.297 0.348 0.291 0.338 0.256 0.325 0.265 0.317 0.263 0.317 0.257 0.316 0.263 0.300 0.264 0.302 0.254 0.297 0.247 0.311 0.246 0.314 0.245 0.306 96 0.141 0.194 0.141 0.194 0.141 0.196 - - - - - - 0.174 0.209 0.159 0.202 0.141 0.175 0.150 0.202 0.149 0.200 0.145 0.195 192 0.196 0.254 0.191 0.249 0.191 0.250 - - - - - - 0.231 0.258 0.210 0.249 0.186 0.220 0.192 0.243 0.194 0.245 0.188 0.240 336 0.260 0.309 0.251 0.302 0.249 0.301 - - - - - - 0.289 0.298 0.265 0.289 0.239 0.262 0.237 0.279 0.236 0.276 0.236 0.277 720 0.346 0.378 0.347 0.380 0.331 0.370 - - - - - - 0.376 0.347 0.362 0.344 0.319 0.315 0.314 0.335 0.304 0.326 0.313 0.335 Weather Avg.
-0.236 0.284 0.233 0.281 0.228 0.279 - - - - - - 0.268 0.278 0.249 0.271 0.221 0.243 0.223 0.265 0.221 0.262 0.221 0.262 • * Weather included in the pre-training of TimesFM is not evaluated, denoted by a dash (–).
-[figure]skip=0pt
+0.236 0.284 0.233 0.281 0.228 0.279 - - - - - - 0.268 0.278 0.249 0.271 0.221 0.243 0.223 0.265 0.221 0.262 0.221 0.262 96 - - - - - - - - - - - - 0.335 0.214 0.333 0.212 0.332 0.211 0.355 0.259 0.355 0.259 0.356 0.259 192 - - - - - - - - - - - - 0.352 0.223 0.351 0.221 0.350 0.220 0.368 0.263 0.369 0.263 0.371 0.265 336 - - - - - - - - - - - - 0.377 0.235 0.375 0.233 0.374 0.231 0.387 0.273 0.385 0.273 0.385 0.273 720 - - - - - - - - - - - - 0.430 0.260 0.426 0.259 0.426 0.257 0.438 0.306 0.431 0.296 0.428 0.294 Traffic Avg.
+- - - - - - - - - - - - 0.374 0.233 0.371 0.231 0.371 0.230 0.387 0.275 0.385 0.273 0.385 0.273 • * The Weather and Traffic datasets included in the pre-training are not evaluated, denoted by a dash (–).
+C.5 Zero-Shot Performance
+We provide comparisons between zero-shot performance and full-shot performance in Table 10.
+Models TimeMoEbase TimeMoElarge TimerXL Moiraibase Moirailarge Chronosbase TTMA PatchTST Methods Zero-shot FT Prune+FT Zero-shot FT Prune+FT Zero-shot FT Prune+FT Zero-shot FT Prune+FT Zero-shot FT Prune+FT Zero-shot FT Prune+FT Zero-shot FT Prune+FT Datasets H MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE 96 0.357 0.381 0.340 0.377 0.340 0.376 0.350 0.382 0.331 0.372 0.329 0.372 0.366 0.392 0.365 0.393 0.353 0.391 0.370 0.397 0.367 0.396 0.366 0.396 0.385 0.395 0.371 0.391 0.370 0.391 0.375 0.375 0.366 0.374 0.363 0.374 0.359 0.394 0.359 0.395 0.358 0.393 0.376 0.396 192 0.384 0.404 0.380 0.408 0.376 0.404 0.388 0.412 0.371 0.398 0.369 0.399 0.401 0.413 0.400 0.416 0.389 0.413 0.404 0.417 0.405 0.418 0.404 0.418 0.427 0.419 0.416 0.416 0.413 0.416 0.432 0.408 0.418 0.405 0.414 0.404 0.389 0.412 0.392 0.417 0.391 0.410 0.399 0.416 336 0.411 0.434 0.406 0.432 0.399 0.425 0.411 0.430 0.392 0.413 0.389 0.411 0.412 0.423 0.419 0.431 0.409 0.426 0.428 0.430 0.435 0.434 0.429 0.430 0.458 0.433 0.455 0.431 0.455 0.432 0.475 0.430 0.453 0.424 0.446 0.422 0.405 0.427 0.407 0.430 0.402 0.424 0.418 0.432 720 0.449 0.477 0.437 0.469 0.424 0.457 0.427 0.455 0.403 0.438 0.387 0.419 0.422 0.442 0.454 0.465 0.433 0.455 0.430 0.442 0.465 0.459 0.427 0.441 0.463 0.448 0.475 0.457 0.454 0.444 0.473 0.447 0.452 0.441 0.446 0.438 0.448 0.466 0.448 0.467 0.449 0.465 0.450 0.469 Avg.
+0.400 0.424 0.391 0.422 0.385 0.416 0.394 0.420 0.374 0.405 0.369 0.400 0.400 0.418 0.410 0.426 0.396 0.421 0.408 0.422 0.418 0.427 0.407 0.421 0.433 0.424 0.429 0.424 0.423 0.421 0.439 0.415 0.422 0.411 0.417 0.410 0.400 0.425 0.402 0.427 0.400 0.423 0.411 0.428 ETTh1 #param 63% 63% 67% 58%69% 36%39% 55% 91%93% 96 0.305 0.359 0.267 0.330 0.262 0.327 0.302 0.354 0.272 0.332 0.272 0.332 0.287 0.345 0.287 0.345 0.275 0.337 0.273 0.338 0.273 0.338 0.272 0.338 0.288 0.341 0.288 0.341 0.287 0.341 0.276 0.315 0.266 0.311 0.266 0.311 0.264 0.333 0.264 0.333 0.262 0.328 0.277 0.339 192 0.351 0.386 0.341 0.380 0.331 0.374 0.364 0.385 0.356 0.383 0.356 0.382 0.344 0.385 0.344 0.385 0.334 0.378 0.332 0.385 0.331 0.385 0.331 0.385 0.352 0.388 0.349 0.386 0.350 0.387 0.348 0.361 0.334 0.357 0.331 0.356 0.321 0.372 0.321 0.372 0.324 0.370 0.345 0.381 336 0.391 0.418 0.386 0.415 0.372 0.403 0.417 0.425 0.420 0.424 0.411 0.418 0.375 0.410 0.375 0.410 0.368 0.404 0.367 0.416 0.358 0.410 0.358 0.409 0.377 0.412 0.373 0.409 0.371 0.408 0.400 0.398 0.378 0.391 0.370 0.388 0.351 0.397 0.351 0.397 0.351 0.396 0.368 0.404 720 0.419 0.454 0.419 0.454 0.377 0.423 0.537 0.496 0.534 0.494 0.459 0.454 0.411 0.445 0.411 0.445 0.398 0.439 0.410 0.456 0.392 0.444 0.388 0.442 0.402 0.439 0.402 0.439 0.394 0.435 0.420 0.421 0.395 0.418 0.389 0.415 0.395 0.439 0.395 0.439 0.393 0.437 0.397 0.432 Avg.
+0.367 0.404 0.353 0.395 0.336 0.382 0.405 0.415 0.396 0.408 0.375 0.397 0.354 0.396 0.354 0.396 0.344 0.390 0.346 0.399 0.339 0.394 0.337 0.394 0.355 0.395 0.353 0.394 0.351 0.393 0.361 0.374 0.343 0.369 0.339 0.368 0.333 0.385 0.333 0.385 0.333 0.383 0.347 0.389 ETTh2 #param 82% 77% 68% 55%78% 27%36% 56% 93%98% 96 0.338 0.368 0.220 0.302 0.220 0.302 0.309 0.357 0.214 0.305 0.211 0.299 0.307 0.355 0.305 0.355 0.297 0.354 0.326 0.360 0.309 0.343 0.304 0.342 0.332 0.356 0.311 0.345 0.312 0.346 0.308 0.323 0.285 0.315 0.282 0.315 0.318 0.348 0.295 0.346 0.291 0.340 0.290 0.343 192 0.353 0.388 0.284 0.347 0.282 0.344 0.346 0.381 0.268 0.345 0.257 0.333 0.350 0.383 0.348 0.383 0.337 0.381 0.349 0.380 0.349 0.367 0.341 0.366 0.355 0.374 0.343 0.370 0.342 0.370 0.365 0.356 0.329 0.345 0.323 0.343 0.354 0.370 0.332 0.366 0.327 0.363 0.329 0.368 336 0.381 0.413 0.349 0.392 0.335 0.381 0.373 0.408 0.326 0.386 0.300 0.364 0.383 0.405 0.382 0.406 0.367 0.403 0.368 0.396 0.379 0.385 0.375 0.384 0.374 0.388 0.364 0.385 0.359 0.385 0.409 0.383 0.361 0.369 0.350 0.365 0.376 0.387 0.356 0.383 0.356 0.381 0.360 0.390 720 0.504 0.493 0.462 0.466 0.415 0.437 0.475 0.477 0.428 0.449 0.380 0.415 0.430 0.438 0.428 0.439 0.409 0.434 0.404 0.418 0.410 0.409 0.397 0.409 0.404 0.411 0.401 0.409 0.400 0.409 0.504 0.430 0.434 0.413 0.412 0.404 0.398 0.406 0.409 0.420 0.395 0.406 0.416 0.422 Avg.
+0.394 0.416 0.329 0.377 0.313 0.366 0.376 0.406 0.309 0.371 0.287 0.353 0.368 0.395 0.366 0.396 0.353 0.393 0.362 0.389 0.362 0.376 0.354 0.375 0.366 0.382 0.355 0.377 0.353 0.378 0.397 0.373 0.352 0.361 0.342 0.357 0.362 0.378 0.348 0.379 0.342 0.373 0.349 0.381 ETTm1 #param 92% 64% 81% 51%60% 51%52% 18% 86%99% 96 0.201 0.291 0.161 0.253 0.158 0.251 0.197 0.286 0.166 0.258 0.164 0.258 0.191 0.279 0.191 0.280 0.181 0.273 0.189 0.275 0.174 0.261 0.172 0.259 0.184 0.268 0.165 0.251 0.165 0.251 0.167 0.236 0.160 0.231 0.156 0.234 0.169 0.256 0.160 0.251 0.160 0.248 0.165 0.254 192 0.258 0.334 0.229 0.308 0.217 0.299 0.250 0.322 0.248 0.318 0.236 0.313 0.245 0.320 0.245 0.320 0.233 0.310 0.242 0.310 0.234 0.304 0.231 0.300 0.246 0.311 0.234 0.304 0.232 0.301 0.238 0.284 0.225 0.276 0.215 0.275 0.223 0.294 0.213 0.289 0.213 0.286 0.221 0.292 336 0.324 0.373 0.313 0.367 0.273 0.342 0.337 0.375 0.347 0.381 0.314 0.367 0.297 0.356 0.297 0.356 0.283 0.344 0.287 0.342 0.284 0.340 0.282 0.337 0.305 0.351 0.288 0.343 0.287 0.340 0.306 0.328 0.286 0.317 0.268 0.313 0.276 0.328 0.269 0.326 0.268 0.322 0.275 0.325 720 0.488 0.464 0.486 0.464 0.375 0.407 0.480 0.461 0.552 0.488 0.471 0.453 0.389 0.416 0.389 0.416 0.367 0.396 0.356 0.388 0.353 0.386 0.350 0.384 0.403 0.413 0.366 0.397 0.365 0.396 0.420 0.395 0.381 0.377 0.348 0.367 0.342 0.373 0.346 0.379 0.338 0.369 0.360 0.380 Avg.
+0.318 0.366 0.297 0.348 0.256 0.325 0.316 0.361 0.328 0.361 0.296 0.348 0.281 0.343 0.281 0.343 0.266 0.331 0.269 0.329 0.261 0.323 0.259 0.320 0.285 0.336 0.263 0.324 0.262 0.322 0.283 0.311 0.263 0.300 0.247 0.297 0.253 0.313 0.247 0.311 0.245 0.306 0.255 0.313 ETTm2 #param 75% 57% 64% 51%56% 35%38% 18% 86%99% 96 0.160 0.214 0.141 0.194 0.141 0.196 0.159 0.213 0.140 0.193 0.139 0.190 0.169 0.225 0.169 0.226 0.165 0.219 0.177 0.243 0.178 0.228 0.159 0.222 0.176 0.248 0.147 0.199 0.147 0.198 0.157 0.191 0.174 0.209 0.143 0.180 0.159 0.211 0.150 0.202 0.145 0.195 0.149 0.196 192 0.210 0.260 0.196 0.254 0.191 0.250 0.215 0.266 0.191 0.245 0.189 0.241 0.217 0.271 0.217 0.271 0.209 0.261 0.220 0.290 0.220 0.269 0.216 0.265 0.219 0.295 0.211 0.263 0.198 0.255 0.204 0.235 0.231 0.258 0.186 0.223 0.203 0.251 0.192 0.243 0.188 0.240 0.193 0.240 336 0.274 0.309 0.260 0.309 0.249 0.301 0.291 0.322 0.255 0.297 0.253 0.292 0.271 0.313 0.267 0.308 0.255 0.296 0.277 0.334 0.308 0.330 0.238 0.287 0.276 0.343 0.277 0.315 0.279 0.317 0.260 0.276 0.289 0.298 0.235 0.262 0.247 0.286 0.237 0.279 0.236 0.277 0.244 0.281 720 0.418 0.405 0.346 0.378 0.331 0.370 0.415 0.400 0.371 0.383 0.361 0.370 0.356 0.373 0.336 0.358 0.322 0.345 0.363 0.392 0.390 0.388 0.333 0.361 0.363 0.407 0.374 0.383 0.345 0.365 0.345 0.331 0.376 0.347 0.296 0.306 0.314 0.335 0.314 0.335 0.313 0.335 0.314 0.332 Avg.
+0.266 0.297 0.236 0.284 0.228 0.279 0.270 0.300 0.239 0.280 0.236 0.273 0.253 0.296 0.247 0.291 0.238 0.280 0.259 0.315 0.274 0.304 0.237 0.284 0.259 0.323 0.252 0.290 0.242 0.284 0.242 0.258 0.268 0.278 0.215 0.243 0.231 0.271 0.223 0.265 0.221 0.262 0.225 0.262 Weather #param 68% 88% 75% 20%62% 20%44% 17% 91%94% 96 0.211 0.343 - - - - - - - - - - 0.141 0.237 0.139 0.237 0.138 0.237 0.215 0.346 0.152 0.247 0.135 0.236 9.889 0.255 0.148 0.243 0.138 0.236 0.121 0.208 0.121 0.208 0.120 0.206 0.152 0.251 0.129 0.226 0.129 0.226 0.133 0.233 192 0.257 0.386 - - - - - - - - - - 0.160 0.255 0.158 0.256 0.157 0.255 0.251 0.382 0.170 0.263 0.162 0.263 0.178 0.255 0.172 0.265 0.164 0.265 0.139 0.226 0.139 0.226 0.138 0.224 0.179 0.275 0.147 0.243 0.148 0.243 0.150 0.248 336 0.281 0.405 - - - - - - - - - - 0.178 0.272 0.177 0.275 0.176 0.275 0.286 0.412 0.192 0.285 0.176 0.275 1.078 0.276 0.187 0.280 0.175 0.283 0.157 0.244 0.157 0.244 0.155 0.242 0.193 0.290 0.167 0.266 0.165 0.263 0.168 0.267 720 0.354 0.465 - - - - - - - - - - 0.216 0.307 0.219 0.313 0.217 0.312 0.355 0.462 0.220 0.307 0.210 0.305 0.268 0.317 0.219 0.309 0.215 0.310 0.199 0.283 0.199 0.283 0.194 0.277 0.243 0.331 0.202 0.297 0.199 0.293 0.202 0.295 Avg.
+0.276 0.400 - - - - - - - - - - 0.174 0.268 0.173 0.270 0.172 0.270 0.277 0.401 0.184 0.276 0.171 0.270 2.853 0.276 0.182 0.274 0.173 0.274 0.154 0.240 0.154 0.240 0.152 0.237 0.192 0.287 0.161 0.258 0.160 0.256 0.163 0.261 ECL #param 56% 40%42% 3%60% 53% 94%98% Win Count 5 0 13 10 19 15 5 0 12 11 14 12 5 1 3 4 9 7 9 5 8 7 11 9 1 5 3 8 5 7 21 27 23 25 29 30 30 13 30 21 30 27 Win Rate 8.3% 46.0% 68.0% 8.3% 46.0% 52.0% 10.0% 11.7% 26.7% 23.3% 25.0% 33.3% 10.0% 18.3% 20.0% 80.0% 80.0% 98.3% 71.7% 85.0% 95.0%
 Appendix D Limitations and Future Works
@@
 As for some tasks where pre-trained TSFMs show poor capabilities and cannot deliver an effective subnetwork, one possible solution is to apply pruning to a fine-tuned TSFM, rewind remaining parameters to the pre-trained weights, and fine-tune the pruned model.
-This follows prior works based on the classic Lottery Ticket Hypothesis [8], while we would like to leave it as a future work.
+This follows prior works based on the classic Lottery Ticket Hypothesis [LotteryTicket], while we would like to leave it as a future work.